### v1

# Dual Consolidation for Pre-Trained Model-Based Domain-Incremental Learning

Abstract
Domain-Incremental Learning (DIL) involves the progressive adaptation of a model to new concepts across different domains. While recent advances in pre-trained models provide a solid foundation for DIL, learning new concepts often results in the catastrophic forgetting of pre-trained knowledge. Specifically, sequential model updates can overwrite both the representation and the classifier with knowledge from the latest domain. Thus, it is crucial to develop a representation and corresponding classifier that accommodate all seen domains throughout the learning process. To this end, we propose DUal ConsolidaTion (Duct) to unify and consolidate historical knowledge at both the representation and classifier levels. By merging the backbone of different stages, we create a representation space suitable for multiple domains incrementally. The merged representation serves as a balanced intermediary that captures task-specific features from all seen domains. Additionally, to address the mismatch between consolidated embeddings and the classifier, we introduce an extra classifier consolidation process. Leveraging class-wise semantic information, we estimate the classifier weights of old domains within the latest embedding space. By merging historical and estimated classifiers, we align them with the consolidated embedding space, facilitating incremental classification. Extensive experimental results on four benchmark datasets demonstrate Duct’s state-of-the-art performance.
1 Introduction
Recent years have seen the rise of deep learning, demonstrating its strong potential in real-world applications He et al. (2016); Deng et al. (2009); Floridi & Chiriatti (2020). However, in a dynamic and ever-changing world, data often evolves in a stream format Aggarwal (2018); Ye et al. (2024), necessitating continuous updates with data from new domains Gama et al. (2014). For instance, autonomous vehicles must handle driving tasks across different seasons and weather conditions Bojarski et al. (2016), and face recognition systems are supposed to recognize users despite the varying lighting conditions Zhao et al. (2003). Correspondingly, Domain-Incremental Learning (DIL) van de Ven et al. (2022) addresses this challenge by absorbing new knowledge from new data distributions while preserving existing knowledge. However, sequentially updating the model with new domains often biases the embedding Yu et al. (2020) and classifier Zhao et al. (2020); Wu et al. (2019) modules toward the latest domain, leading to catastrophic forgetting Mermillod et al. (2013); De Lange et al. (2021) of previous knowledge. Therefore, developing methods for learning new domains without forgetting existing knowledge becomes a critical challenge for the machine learning community.
To combat catastrophic forgetting, recent works leverage strong pre-trained models (PTMs)Han et al. (2021) as the initialization for DILWang et al. (2022d; c; b). PTMs’ robust representation abilities provide a powerful starting point, showing promising results in DIL benchmarks. These approaches freeze the pre-trained backbone to preserve existing knowledge and append lightweight modules Jia et al. (2022); Chen et al. (2022); Lian et al. (2022) to capture new patterns. However, since incoming domains will also overwrite the appended modules, the representations still suffer from forgetting, leading to corresponding bias in the classifier.
In DIL, there are two primary sources of forgetting: the overwriting of features and classifiers. Continually adjusting the features to capture the latest task biases the representations toward the newest domain. For example, suppose the previous task involves real-world photos while the subsequent task includes quick-draw images. The features will be overwhelmingly adjusted to highlight non-objective details, which are less helpful in distinguishing the previous domain, thus leading to forgetting. In light of this, an ideal feature space in DIL should fit all domains and equally highlight the domain-specific features for all seen tasks. However, as features become less generalizable due to overwriting, the classifier also becomes biased toward the latest domain. Since the classifier plays the pivotal role of mapping embeddings to classes, using the biased classifier can cause detrimental effects to decision-making. Consequently, it is imperative to consolidate existing knowledge both in the feature and classifier to resist forgetting.
There are two main challenges to achieving this goal. 1) Building a unified embedding space that suits all tasks continually. Since the training data arrives in a stream format, we cannot access all training instances simultaneously, preventing us from achieving the ‘oracle embedding’ that favors all seen domains. 2) Calibrating the prediction of biased classifiers. Even if we achieve a holistic embedding suitable for all tasks, we still cannot build an accurate mapping between features and classes due to the lack of previous instances. Consequently, a classifier consolidation process is needed to align the classifier weights with the changing features continually.
In this paper, we propose DUal ConsolidaTion (Duct) to address these challenges. To counter feature drift, we introduce a representation merging technique that continuously integrates historical backbones. This merged embedding consolidates task-specific information from all previous domains, providing informative features without forgetting. Consequently, we achieve non-forgettable features incrementally and resist feature-level forgetting. To map updated features to classes, we design a classifier consolidation process that merges historical classifier weights with calibrated weights. By extracting class-wise semantic relationships in the joint space, we develop a classifier transport mechanism that estimates classifiers of previous domain classes in the latest embedding space. Through consolidating features and classifiers in this coordinated way, Duct balances all seen domains and robustly resists forgetting, achieving competitive performance on various benchmarks.
2 Related Work
Continual Learning/Incremental Learning is a popular topic in the machine learning field, aiming to incorporate new knowledge within streaming data De Lange et al. (2021); Masana et al. (2023); Wang et al. (2024); Zhao et al. (2021); Zheng et al. (2024); Qi et al. (2024). It can be further classified into three subcategories, i.e., task-incremental learning (TIL) De Lange et al. (2021), class-incremental learning (CIL) Masana et al. (2023), and domain-incremental learning (DIL) van de Ven et al. (2022). Among them, TIL and CIL are faced with emerging new classes and are required to learn them without forgetting. By contrast, DIL mainly focuses on the scenario where label space is fixed while incoming data involves new domains Wang et al. (2022b); Shi & Wang (2023). In continual learning, typical solutions include using data replay Bang et al. (2021); Aljundi et al. (2019b); Isele & Cosgun (2018); Rolnick et al. (2019); Ratcliff (1990) to review former knowledge, using knowledge distillation Hinton et al. (2015); Rebuffi et al. (2017); Li & Hoiem (2017) or parameter regularization Kirkpatrick et al. (2017); Zenke et al. (2017); Chaudhry et al. (2018); Aljundi et al. (2018; 2019a) to maintain existing knowledge. Recent advances also resort to bias correction Hou et al. (2019); Zhao et al. (2020); Ahn et al. (2021); Castro et al. (2018); Wu et al. (2019); Belouadah & Popescu (2019) to rectify the model bias or expand the network structure as data evolves Rusu et al. (2016); Aljundi et al. (2017); Yan et al. (2021); Wang et al. (2022a); Zhou et al. (2023).
Domain-Incremental Learning with Pre-Trained Models: With the rapid development of pre-training techniques, PTMs provide a strong initialization for DIL models to boost the feature generalizability Smith et al. (2023); Wang et al. (2022c; d; b). Current PTM-based DIL methods mainly resort to visual prompt tuning Jia et al. (2022) to adjust the pre-trained features. With the pre-trained features frozen, it learns a prompt pool as external knowledge and selects instance-specific prompts to encode task information. L2P Wang et al. (2022d) utilizes the query-key matching mechanism to select instance-specific prompts, while DualPrompt Wang et al. (2022c) extends it by learning task-specific and instance-specific prompts jointly. By contrast, S-Prompts Wang et al. (2022b) learns domain-specific prompts for DIL and retrieves prompts via KNN search. To alleviate the prompt selection cost, CODA-Prompt Smith et al. (2023) modifies the prompt selection process into an attention-based weighted combination. Apart from prompt tuning, there are also some methods that directly build classifiers upon the pre-trained features Zhou et al. (2024a); McDonnell et al. (2023).
Model Merging: Recent advances in pre-training have made weight interpolation among different models a useful technique in model editing Ainsworth et al. (2023); Frankle et al. (2020); Ilharco et al. (2023); Matena & Raffel (2021); Wortsman et al. (2022a; b). Most work focuses on simple averaging of multiple models, aiming to enhance model in terms of its performance on the single task and robustness to out-of-distribution data. In this paper, we consider merging the pre-trained weights and task vector parameters to cultivate multitask representation for all seen domains.
3 From Old Domains to New Domains
In this section, we introduce the background information about domain-incremental learning and the baselines for applying pre-trained models to domain-incremental learning.
3.1 Domain-Incremental Learning
In domain-incremental learning, the model is faced with the data stream with a sequence of tasks $\left\{\mathcal{D}^{1},\mathcal{D}^{2},\cdots,\mathcal{D}^{B}\right\}$, where $\mathcal{D}^{b}=\left\{\mathcal{X}_{b},\mathcal{Y}_{b}\right\}$ is the $b$-th training task, i.e., $\mathcal{X}_{b}=\left\{{\bf x}_{i}\right\}_{i=1}^{n_{b}}$ and $\mathcal{Y}_{b}=\left\{y_{i}\right\}_{i=1}^{n_{b}}$. Each training instance ${\bf x}_{i}\in\mathbb{R}^{D}$ belongs to class $y_{i}\in Y$. This paper follows the exemplar-free setting Zhu et al. (2021); Wang et al. (2022d), i.e., during the $b$-th training stage, we can only access data in the current dataset $\mathcal{D}^{b}$. In DIL, the label space $Y$ is unchanged throughout the learning process, while the distribution of input instance keeps changing from domain to domain, i.e., $p(\mathcal{X}_{b})\neq p(\mathcal{X}_{b^{\prime}})$ for $b\neq b^{\prime}$. The target of DIL is to fit a model $f({\bf x})$ that can discriminate the classes among any seen domains:
$f^{*}=\underset{f\in\mathcal{H}}{\operatorname{argmin}}\enspace\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{t}^{1}\cup\cdots\mathcal{D}_{t}^{b}}\mathbb{I}(y\neq f(\mathbf{x}))\,,$ (1)
where $\mathcal{H}$ is the hypothesis space, $\mathbb{I}(\cdot)$ is the indicator function which outputs $1$ if the expression holds and 00 otherwise. $\mathcal{D}_{t}^{b}$ denotes the data distribution of task $b$. Consequently, DIL models are supposed to classify instances of new domains while not forgetting previous ones.
Following Wang et al. (2022c; d; b); Smith et al. (2023), we assume a pre-trained Vision Transformer (ViT) Dosovitskiy et al. (2020) is available as the initialization for $f({\bf x})$. For simplicity, we decouple the network structure into the embedding function $\phi(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}^{d}$ (i.e., the final [CLS] token) and the linear classifier $W\in\mathbb{R}^{d\times|Y|}$. In this way, the output is denoted as $f({\bf x})=W^{\top}\phi({\bf x})$, and we utilize a cosine classifier in this paper. The classifier is further decoupled into $W=[{\bf w}_{1},{\bf w}_{2},\cdots,{\bf w}_{|{Y}|}]$, where ${\bf w}_{j}$ denotes the classifier weight of the $j$-th class. Although the label space is static during the learning process, data distribution of the same class across different domains can yield significant domain gaps. Hence, we follow existing works Wang et al. (2022c; d; b); Smith et al. (2023) to expand the classifier $W\in\mathbb{R}^{d\times b|Y|}$ as new tasks emerge, and the final prediction is made by $\left(\operatorname*{arg\,max}_{i}{\bf w}_{i}^{\top}\phi({\bf x})\right)\textbf{ mod }|Y|$.
3.2 Baselines in Domain-Incremental Learning
In domain-incremental learning, a naive solution facing the incoming datasets of new domains is to directly optimize the embedding and classifier:
$\min_{W\cup\phi}\sum_{({\bf x},y)\in{D^{b}}}\ell\left(W^{\top}\phi({\bf x}),y\right)\,.$ (2)
With pre-trained weights as initialization, Eq. 2 can quickly capture the discriminative features within the new domain. However, since the embedding $\phi(\cdot)$ is kept changing among different domains, it quickly loses the generalization ability to previous domains and suffers from forgetting.
To resist feature-level forgetting, a feasible solution is to freeze the pre-trained weights and append lightweight modules to encode domain-specific knowledge, e.g., prompt Jia et al. (2022). A representative work L2P Wang et al. (2022d) organizes a set of prompts as the prompt pool (i.e., $\mathbf{Pool}$) and optimizes the model via:
$\min_{\mathbf{Pool}\cup W}\sum_{({\bf x},y)\in{D^{b}}}\ell\left(W^{\top}\bar{\phi}\left({\bf x};\mathbf{Pool}\right),y\right)+\mathcal{L}_{\mathbf{Pool}}\,,$ (3)
where $\bar{\phi}\left({\bf x};\mathbf{Pool}\right)$ denotes the prompted feature representation with backbone frozen, and $\mathcal{L}_{\mathbf{Pool}}$ denotes the prompt selection loss Wang et al. (2022d). Eq. 3 dynamically retrieves instance-specific prompts to adjust the representations and learns the classifier to map the features to corresponding classes. Since the backbone weights are frozen, it can alleviate the representation drift during updating.
Discussions: Although Eq. 2 and Eq. 3 adopt different policies for model updating, both of them suffer from the feature-level forgetting. Specifically, tuning the model via Eq. 2 can encode task-specific knowledge into the embedding, but also comes with the risk of feature being fully overwritten by new domains since there is no restriction on retaining previous knowledge. Besides, although Eq. 3 freezes the pre-trained weights, the trainable prompts appended may also incur forgetting. Since the prompted feature relies on instance-specific prompt selection, optimizing the prompts for new domains still has the risk of overwriting previous ones. With the features being biased toward the latest domain, the classifier is also adjusted upon it. Consequently, the classifiers of previous tasks are incompatible with the unstable and ever-changing features, resulting in the classifier-level forgetting.
4 Duct: Dual Consolidation for Domain-Incremental Learning
Observing the risk of feature-level and classifier-level overwriting, we aim to resist forgetting in DIL from two aspects. Firstly, alleviating feature-level forgetting requires obtaining a unified embedding space at a low cost. Since sequentially updating the backbone or prompts will both result in forgetting, a proper solution is needed to make full use of all historical features and consolidate them into a unified one. Secondly, as the representation changes from task to task, the mismatch between classifier and embedding becomes more severe as the data evolves. Consequently, a classifier consolidation step is needed to calibrate them to be compatible with the embeddings.
In the following sections, we introduce the learning paradigm for representation consolidation and classifier consolidation. Lastly, we provide detailed guidelines for training and inference.
4.1 Representation Consolidation
Observing the challenge of striking a balance between all seen domains, we need to opt for another feature updating policy to avoid sequential overwriting and resist forgetting. Let us assume an ideal scenario where we can separately train each model for each incoming domain, obtaining a set of models $\{\phi_{1}(\cdot),W_{1}\},\cdots,\{\phi_{B}(\cdot),W_{B}\}$. Those models are obtained by optimizing the same PTM via Eq. 2, thus having the discriminability for each specific domain and can be seen as the ‘domain expert.’ In an oversimplified scenario where we know which domain the instance is from, we can directly utilize the corresponding expert for prediction. However, since such auxiliary information is unavailable in DIL, we need to obtain a omnipotent embedding space that suits all domains. In this way, there is no need to decide which embedding to use since the omnipotent embedding is strong enough to capture all domain-specific features.
Correspondingly, we take inspiration from the model merging community Ilharco et al. (2023); Ramé et al. (2023) that an ideal model for multiple domains can be achieved by combining multiple task vectors. We denote the relative change of the embedding function as $\delta_{\phi_{i}}=\phi_{i}-\phi_{0}$, where $\phi_{0}$ is the weight of the pre-trained model. Hence, $\delta_{\phi_{i}}$ represents the relative change of the weights in the $i$-th domain, and we can build the unified embedding via:
$\phi^{m}_{i}=\phi_{0}+\alpha_{\phi}\sum_{i}\delta_{\phi_{i}}\,,$ (4)
which combines the pre-trained weight with every task vector. Since all the fine-tuned embeddings start from the same pre-trained weight, Ilharco et al. (2023); Zhang & Bottou (2023) verify the task vectors share low similarity due to the semantic gap between different domains. Consequently, adding such weights enables the merged model to highlight all task-specific features, and Eq. 4 provides a feasible way to obtain the universal feature representation that suits all seen domains.
Representation consolidation with task similarity: Although Eq. 4 provides a simple way to consolidate multiple task vectors to build a unified embedding, it still lacks consideration in measuring task-wise similarities. For example, if two distinct domains are more similar, highlighting those features would be more helpful for recognizing classes of corresponding domains. Hence, we introduce task-similarity ($\textbf{Sim}_{0,i}$) into the representation consolidation process and replace Eq. 4 into:
$\phi^{m}_{i}=\phi_{0}+\alpha_{\phi}\sum_{i}\textbf{Sim}_{0,i}\delta_{\phi_{i}}\,,$ (5)
Specifically, since the merged embedding is built upon pre-trained weights and task vectors, we consider the similarity of pre-trained model $\phi_{0}$ and subsequent models $\phi_{i}$ to adjust the merging process. A naive way to is to directly measure the cosine similarity between $\phi_{0}$ and $\phi_{i}$. However, since pre-trained weights are with millions of dimensions, the similarity in such space is ineffective due to curse of dimensionality. To this end, we utilize the current training task as an indicator for task similarity. Specifically, for every class in $\mathcal{D}^{b}$, we utilize the backbone $\phi_{i}$ to extract class centers in its embedding space:
${\bm{c}}^{i}_{p}=\sum_{j=1}^{|\mathcal{{D}}^{b}|}\mathbb{I}(y_{j}=p)\phi_{i}({\bf x}_{j})/\sum_{j=1}^{|\mathcal{{D}}^{b}|}\mathbb{I}(y_{j}=p)\,.$ (6)
Afterward, we can obtain two sets of class centers in the embedding space, i.e., ${\bm{C}^{0}}=[{\bm{c}}^{0}_{1},{\bm{c}}^{0}_{2},\cdots,{\bm{c}}^{0}_{|{Y}|}]$, ${\bm{C}^{i}}=[{\bm{c}}^{i}_{1},{\bm{c}}^{i}_{2},\cdots,{\bm{c}}^{i}_{|{Y}|}]$. Since those class centers are representative points in the embedding space, we calculate the pair-wise class similarity to indicate task similarities:
$\textbf{Sim}_{0,i}=\frac{1}{|{Y}|}\sum^{|{Y}|}_{j=1}\text{sim}({\bm{c}}^{0}_{j},{\bm{c}}^{i}_{j})\,,$ (7)
where we utilize cosine similarity to calculate center-wise similarity $\text{sim}(\cdot,\cdot)$.
Effect of representation consolidation: Figure 1(top) visualizes the representation consolidation process, where the merged backbone can unify multiple domains. Through Eq. 5, we are able to build the joint embedding space that suits all tasks by combining them in the parameter space. Since tasks in DIL emerge one-by-one, the representation consolidation process can also be made incrementally, i.e., $\phi^{m}_{i}=\phi^{m}_{i-1}+\alpha_{\phi}\textbf{Sim}_{0,i}\delta_{\phi_{i}}$. In this way, we can get rid of the extensive memory cost and only keep at most two backbones in memory. In each training stage, we first fine-tune the model via Eq. 2 and conduct representation consolidation to aggregate the representations. In this way, we obtain a unified embedding space across all seen tasks.
Figure 1: Illustration of Duct. Top: Representation consolidation. We utilize the pre-trained model as initialization and optimize it for each domain, obtaining the task vectors. Afterward, we combine the pre-trained model and all seen task vectors to build the unified embedding space. Bottom: Classifier consolidation. To align the classifiers with consolidated features, we design the new classifier retraining and old classifier transport to consolidate classifiers. Class-wise semantic information is utilized in classifier transport.
4.2 Classifier Consolidation
In Eq. 5, we design the feature consolidation process, which aggregates multiple fine-tuned models by accumulating task vectors. However, a fatal problem still exists, i.e., there is a mismatch between the classifiers and the consolidated embeddings. Since the classifiers are optimized to match the embeddings to the corresponding class, the matching degree drastically decays as the backbone is replaced with another one. Recalling the background information in Section 3.1 that we utilize an independent classifier for each domain, we need to design the extra classifier consolidation process to align the classifiers to the embedding space. In each training stage, we denote the classifier for previous domains as ‘old classifier’ ($W_{o}$) and the classifier for the current domain as ‘new classifier’ ($W_{n}$). We then discuss how to align them with the consolidated features.
New Classifier Retraining: In each training stage, we have the training data $\mathcal{D}^{b}$ in hand. Consequently, it is intuitive to coordinate the new classifier with the consolidated features via:
$\min_{W_{n}}\sum_{({\bf x},y)\in{D^{b}}}\ell\left(W_{n}^{\top}\bar{\phi}_{i}^{m}({\bf x}),y\right)\,,$ (8)
where the consolidated feature $\bar{\phi}_{i}^{m}(\cdot)$ is frozen to resist further mismatch, and we can align the new classifier ${W_{n}}$ to match the features with Eq. 8.
Old Classifier Transport: Eq. 8 aligns the new classifier to the latest embedding space. However, a fatal problem still exists, i.e., the old classifier is also incompatible with the merged embedding space. Consequently, the previous knowledge of old domains shall be forgotten in the incremental learning process. If we have plenty of training instances of previous domains, a similar calibration process can be done as in Eq. 8. However, due to the exemplar-free restrictions, we cannot save any previous instances in the memory. Hence, we need to find another way to estimate the relative calibration weight for the old classifier. For example, if we have the retrained classifier to classify a ‘lion’ in the clip art style, it can be mostly reused to classify the lion in the real photo style. We call such a relationship ‘semantic information,’ and the goal is to utilize such information to assist old classifier alignment. We denote the calibrated classifier using semantic information as $\hat{W}_{o}=\mathcal{T}(W_{n},\textbf{S})$, i.e., the estimated classifier is obtained by transforming the new classifier using semantic information S. Hence, there remain two core problems to be solved: 1) How to define the transformation function $\mathcal{T}$? 2) How to define the semantic information S?
Implementing $\mathcal{T}$ via Optimal Transport: The function $\mathcal{T}$ encodes the correlation between the set of features among two tasks. Considering that the weight matrix of a linear layer reveals the relative relationship among feature-class pairs and the final logit is the aggregation of all features, we can design the recombination of exiting classifiers with a linear mapping $T\in\mathbb{R}^{\beta\times\gamma}$. $T$ encodes the cross-task correlation between the current and the previous domains, where larger values in $T$ denote higher class-wise similarity. Since the decision is made by matching the classifier with corresponding features, we can utilize and recombine the weights of similar classes in the current domain to obtain those of previous domains. For example, important features that discriminate lions in the clip art style should be assigned higher coefficients to help classify lions in the photo style and vice versa.
We denote $\bm{\mu}_{1}\in\Delta_{\beta}$ and $\bm{\mu}_{2}\in\Delta_{\gamma}$, where $\Delta_{d}=\left\{\bm{\mu}:\bm{\mu}\in\mathbb{R}_{+}^{d},\bm{\mu}^{\top}\mathbf{1}=1\right\}$ is the $d$-dimensional simplex. $\bm{\mu}_{1}$ and $\bm{\mu}_{2}$ denote the importance of each class across domains, and we set them as uniform distribution. To map a set of classes to another between old and new domains, a cost matrix $Q\in\mathbb{R}_{+}^{\beta\times\gamma}$ is further introduced to guide the transition. The larger weight of $Q_{i,j}$ indicates we need to pay more cost when reusing the classifier of $i$-th class to assist the $j$-th class. Consequently, the matrix $T$ can be formulated as the coupling of two distributions, aiming to connect classes from different domains at the lowest transportation cost. Hence, $T$ can be obtained via minimizing:
$\min_{T}\langle T,Q\rangle\quad\text{ s.t. }T\mathbf{1}=\bm{\mu}_{1},T^{\top}\mathbf{1}=\bm{\mu}_{2},T\geq 0\,,$ (9)
which is the Kantorovich formulation of optimal transport (OT) Villani (2008); Peyré et al. (2019). Optimizing Eq. 9 enables us to show how to move the probability mass across domains with low cost. Since the target is to reuse the retrained classifier $W_{n}$ to adjust previous classifier $W_{o}$, we set $\bm{\mu}_{1}\in\Delta_{|Y|}$ and $\bm{\mu}_{2}\in\Delta_{|Y_{o}|}$ with uniform class marginals, where $|Y_{o}|$ denotes the number of classes in $W_{o}$. Solving Eq. 9 produces the alignment between different domains, and we can apply $T$ to the new classifier to obtain the estimated old classifier, i.e., $\hat{W}_{o}=W_{n}T$.
Defining the Transportation Cost with Semantic Information: Solving Eq. 9 requires a proper definition of the cross-domain cost, i.e., $Q$. The higher cost indicates it is less effective to transport the classifier to the target class and vice versa. To this end, we design a simple yet effective way to measure such class-wise information. Specifically, before the training of each stage, we utilize the pre-trained backbone $\phi_{0}$ to extract class centers in its embedding space via Eq. 6, i.e., $[{\bm{c}}^{0}_{1},{\bm{c}}^{0}_{2},\cdots,{\bm{c}}^{0}_{|{Y}|}]$. These class centers indicate the most representative embedding of the corresponding classes. Since $\phi_{0}$ is optimized with extensive datasets, we rely on its generalizability to reflect task-wise similarity. If two classes are similar, their embeddings should also be situated near each other. Consequently, we calculate the Euclidean distance between class centers as the transportation cost, i.e., $Q_{i,j}=\left\|{\bm{c}}^{0}_{i}-{\bm{c}}^{0}_{j}\right\|_{2}^{2}$. Here classes $i$ and $j$ are from different domains. With the pair-wise transportation cost, we are able to optimize Eq. 9 for the transportation plan $T$. Afterward, we utilize the interpolation between old classifiers and the transported classifier for consolidation:
$W_{o}^{m}=(1-\alpha_{W})W_{o}+\alpha_{W}\hat{W}_{o}=(1-\alpha_{W})W_{o}+\alpha_{W}W_{n}T\,.$ (10)
Effect of classifier consolidation: We visualize the classifier consolidation process in Figure 1 (bottom). The classifier consolidation process takes two steps, i.e., new classifier retraining and old classifier transport. The first step is designed to align the new classifiers with the unified embedding, while the second step aims to recombine the old classifiers with the calibrated classifier. In this way, we obtain a compatible classifier with the consolidated features, thus favoring the recognition in the unified embedding space.
4.3 Summary of Duct
Algorithm 1 Duct for DIL
Input: Incremental datasets: $\left\{\mathcal{D}^{1},\mathcal{D}^{2},\cdots,\mathcal{D}^{B}\right\}$, Pre-trained embedding: $\phi_{0}({\bf x})$;
Output: Incrementally trained model;
1:for $b=1,2\cdots,B$ do
2:     Get the incremental training set $\mathcal{D}^{b}$;
3:     Extract class centers via Eq. 6;
4:     Optimize the model via Eq. 2 ;
5:     Consolidate representations via Eq. 5;
6:     Retrain new classifiers via Eq. 8;
7:     Solve OT via Eq. 9;
8:     Consolidate old classifiers via Eq. 10; return the updated model;
We summarize the training steps in Algorithm 1. Specifically, facing a new training task, we extract the class centers with $\phi_{0}$. Afterward, we optimize the model and separately consolidate the features and classifiers. The consolidated model ($[W_{o}^{m};W_{n}]^{\top}\phi^{m}_{i}$) is then utilized for testing. We utilize Sinkhorn’s algorithm Sinkhorn & Knopp (1967) to solve the OT problem in Eq. 9.
Discussions on memory cost: During the training process, we need to keep two models in the memory, i.e., the historical consolidated backbone $\phi^{m}_{i}$ and the online backbone $\phi_{i}$ for learning the current task. However, after the learning process of each domain, the current running model is merged into $\phi^{m}_{i}$ (Eq. 5) without needing to be kept in memory. During inference, we utilize the consolidated backbone $\phi^{m}_{i}$ and classifiers $[W_{o}^{m};W_{n}]$ for classification, which is the same as a single backbone. Consequently, Duct can be fairly compared with others.
5 Experiment
In this section, we conduct experiments on benchmark datasets to compare Duct to existing state-of-the-art methods. We also provide ablation studies on the components and parameter robustness to analyze the effect of different modules. Visualizations of the embedding space also verify the effectiveness of our proposed method.
(a) Office-Home ViT-B/16 IN1K
(b) DomainNet ViT-B/16 IN1K
(c) CORe50 ViT-B/16 IN1K
(d) CDDB ViT-B/16 IN1K
(e) DomainNet ViT-B/16 IN21K
(f) CORe50 ViT-B/16 IN21K
Figure 2: Incremental performance of different methods with the same pre-trained model. We report the performance gap after the last incremental stage between Duct and the runner-up method at the end of the line.
Table 1: Average and last performance of different methods among five task orders. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implemented with exemplars (10 per class).
Method Office-Home DomainNet CORe50 CDDB
$\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$
Finetune 78.32$\pm$3.28 76.16$\pm$1.39 28.17$\pm$6.47 38.82$\pm$7.65 75.44$\pm$1.68 76.19$\pm$2.36 52.08$\pm$1.35 50.11$\pm$1.62
Replay Ratcliff (1990) 84.23$\pm$2.31 83.75$\pm$0.68 64.78$\pm$2.98 61.16$\pm$1.19 85.56$\pm$0.38 92.21$\pm$0.63 66.91$\pm$18.0 63.21$\pm$11.6
iCaRL Rebuffi et al. (2017) 81.66$\pm$2.43 81.11$\pm$1.23 59.89$\pm$2.86 57.46$\pm$2.31 74.43$\pm$3.18 79.86$\pm$2.96 68.43$\pm$18.7 70.50$\pm$16.5
MEMO Zhou et al. (2023) 71.18$\pm$2.76 63.09$\pm$1.80 61.92$\pm$5.39 58.41$\pm$3.20 64.80$\pm$3.16 68.24$\pm$2.41 60.87$\pm$13.4 58.09$\pm$11.7
SimpleCIL Zhou et al. (2024a) 75.69$\pm$5.03 75.72$\pm$0.00 42.95$\pm$4.84 44.08$\pm$0.00 70.92$\pm$0.74 74.80$\pm$0.00 60.80$\pm$4.49 63.40$\pm$0.00
L2P Wang et al. (2022d) 79.72$\pm$4.19 80.03$\pm$1.29 50.45$\pm$4.10 48.72$\pm$2.83 83.57$\pm$0.35 87.87$\pm$0.51 67.33$\pm$5.98 64.45$\pm$5.83
DualPrompt Wang et al. (2022c) 80.20$\pm$3.81 80.85$\pm$0.14 52.28$\pm$3.35 50.46$\pm$3.17 84.53$\pm$0.89 87.27$\pm$1.06 68.33$\pm$7.52 71.41$\pm$1.24
CODA-Prompt Smith et al. (2023) 84.70$\pm$2.94 85.07$\pm$0.34 59.85$\pm$4.49 59.99$\pm$0.88 87.92$\pm$0.41 91.57$\pm$0.69 69.19$\pm$6.11 74.18$\pm$1.33
EASE Zhou et al. (2024b) 81.16$\pm$3.52 76.33$\pm$2.16 50.50$\pm$2.27 43.72$\pm$1.70 86.30$\pm$0.04 87.02$\pm$1.21 67.78$\pm$2.44 64.96$\pm$8.36
RanPAC McDonnell et al. (2023) 82.30$\pm$3.34 82.28$\pm$0.00 55.20$\pm$3.93 54.80$\pm$0.36 79.16$\pm$0.55 81.38$\pm$0.12 78.92$\pm$4.85 80.48$\pm$0.68
S-iPrompt Wang et al. (2022b) 81.50$\pm$3.17 80.51$\pm$0.21 61.16$\pm$4.57 60.46$\pm$0.90 81.95$\pm$0.57 83.38$\pm$0.70 68.51$\pm$7.20 72.76$\pm$0.43
Duct 86.27$\pm$2.95 86.91$\pm$0.06 67.16$\pm$3.75 67.01$\pm$1.35 91.95$\pm$0.15 94.47$\pm$0.33 84.14$\pm$4.37 85.10$\pm$0.52
5.1 Experimental Setup
Dataset: Following the benchmark settings Wang et al. (2022c; d; b); Smith et al. (2023) in PTM-based domain-incremental learning, we evaluate the performance on Office-Home Venkateswara et al. (2017), DomainNet Peng et al. (2019), CORe50 Lomonaco & Maltoni (2017), and CDDB-Hard Li et al. (2023). Specifically, Office-Home has four domains, DomainNet has six domains, CORe50 has 11 domains, and CDDB-Hard has five domains. We consider five task orders to shuffle these domains in the DIL setting for a holistic evaluation and report the details in Section B.2.
Comparison methods: In the comparison, we consider two types of methods, including exemplar-based methods, i.e., Replay Ratcliff (1990), iCaRL Rebuffi et al. (2017), MEMO Zhou et al. (2023), and SOTA exemplar-free DIL methods, i.e., SimpleCIL Zhou et al. (2024a), L2P Wang et al. (2022d), DualPrompt Wang et al. (2022c), CODA-Prompt Smith et al. (2023), EASE Zhou et al. (2024b), RanPAC McDonnell et al. (2023), and S-iPrompt Wang et al. (2022b). We utilize the same pre-trained backbone for all compared methods.
Implementation details: We deploy the experiments using PyTorch Paszke et al. (2019) and Pilot Sun et al. (2023) on NVIDIA 4090. Following Wang et al. (2022d); Zhou et al. (2024b), we consider two typical pre-trained weights, i.e., ViT-B/16-IN21K and ViT-B/16-IN1K. Both are pre-trained with ImageNet21K Russakovsky et al. (2015), while the latter is further finetuned on ImageNet1K. We optimize Duct using SGD optimizer with a batch size of 128 for 15 epochs. The learning rate is set to $0.001$. We select 10 exemplars per class for exemplar-based methods using herding Welling (2009) algorithm. In Duct, we set the consolidation parameter $\alpha_{\phi}=0.5,\alpha_{w}=0.5$. The code will be made publicly available upon acceptance.
Performance Measure: Following Wang et al. (2022d; b), we denote the accuracy among all seen domains after learning the $b$-th task as $\mathcal{A}_{b}$, we mainly consider $\mathcal{A}_{B}$ (the performance after learning the last stage) and $\bar{\mathcal{A}}=\frac{1}{B}\sum_{b=1}^{B}\mathcal{A}_{b}$ (the average performance among all incremental stages) for comparison. We also consider the forgetting measure Chaudhry et al. (2018) to measure the relative forgetting degree in DIL.
5.2 Benchmark Comparison
[…]
Besides, we also report the incremental performance among five task orders (average and standard deviation) in Table 1 and Figure 3(a). As we can infer from the table, Duct works robustly among different task orders, showing stable improvements against other state-of-the-art methods.
(a) Multiple trails
(b) Forgetting measure $\downarrow$
(c) Parameter robustness
Figure 3: Further analysis on multiple task orders, forgetting measure, and parameter robustness. (a): Incremental performance of different methods on CORe50 with five task orders. The shadow indicates standard deviation. (b): Forgetting measure (lower is better) of different methods on CDDB dataset among five task orders. Duct shows the least forgetting among all methods. (c): Average incremental performance with change of the consolidation ratios.
5.3 Further Analysis
Forgetting Measure: Figure 2 and Table 1 mainly focus on the accuracy measure, which utilizes all seen domains to measure the relative performance. Apart from the accuracy measure, we also follow Wang et al. (2022b; d) to utilize the forgetting measure, which captures the stability of existing knowledge among previous domains. As shown in Figure 3(b), we report the forgetting measure among all compared methods on the CDDB dataset. As we can infer from the figure, several methods suffer from severe forgetting, e.g., Finetune and MEMO, indicating that they are unsuitable for the domain-incremental learning scenario. We also observe that prompt-based methods freeze the backbone to resist feature drift, which tackles the forgetting phenomena. However, Duct still shows the least forgetting (i.e., 0.12) among all compared methods, indicating its strong performance.
Parameter Robustness: There are two major consolidation steps in Duct, i.e., the representation consolidation in Eq. 5 and the classifier consolidation in Eq. 10. These consolidation steps include the merging parameters between a set of models/classifiers, i.e., the backbone merge ratio $\alpha_{\phi}$ and the classifier merge ratio $\alpha_{W}$. In this section, we conduct ablations on the choice of these parameters to investigate the robustness of Duct with change of them. Specifically, we choose $\alpha_{\phi}$ and $\alpha_{W}$ among $\{0.1,0.25,0.5,0.75,1.0\}$, resulting in 25 parameter combinations. We conduct experiments with these parameter combinations on the Office-Home dataset and show the average performance in Figure 3(c). As we can infer from the figure, Duct is generally robust to parameter changes. Besides, since the merge parameters are designed to trade-off between old and new knowledge, we find a small value (e.g., 0.1) or large value (e.g., 1.0) works poorly. The poor performance of these marginal parameters also indicates the importance of the merging process since a small value or large value equals to ignoring part of the important information during consolidation. By contrast, both of them prefer a mild value, i.e., choosing them around {0.25, 0.5} shows better performance. Hence, we suggest setting $\alpha_{\phi}=0.5,\alpha_{w}=0.5$ as the default setting.
Variations $\bar{\mathcal{A}}$ ${\mathcal{A}}_{B}$
Baseline 66.56 63.40
Variation 1 80.36 74.17
Variation 2 81.41 76.82
Variation 3 85.42 80.31
Duct 87.74 82.35
Table 2: Ablation study on differnt modules in Duct.
Figure 4: Before Duct.
Figure 5: After Duct.
Figure 6: Visualizations of the embedding space via t-SNE Van der Maaten & Hinton (2008) on DomainNet. We visualize two incremental stages and utilize dots to represent the first domain and triangles for the second domain. The shadow region denotes the decision boundaries.
Compositional Components: In this section, we conduct experiments to investigate the importance of each module in Duct on CDDB dataset. As shown in Table 2, ‘Baseline’ denotes directly freezing the embedding and extracting class centers as the classifier, which shows poor performance due to the domain gap to the downstream domains. We then utilize Eq. 4 to sequentially merge the backbone to consolidate the representations and denote it as Variation 1. It shows that adding the representation consolidation process drastically improves the performance, indicating the superiority of such unified representation in DIL. However, when switching Eq. 4 to Eq. 5, we find Variation 2 further improves the performance, indicating that task similarity is helpful in obtaining a universal embedding space. Afterward, we combine it with the classifier retraining process in Eq. 8, which is denoted as Variation 3. It shows that the mismatch between consolidated features and the classifiers weakens the performance, and aligning the new classifiers to the embedding helps DIL. When comparing Duct to Variation 3, we find the old classifier transport is also vital to recover former knowledge and resist forgetting, which improves the final accuracy by 2%. Ablation studies verify the efficacy of different modules in Duct.
Visualizations: In this section, we visualize the embedding space to show the effectiveness of Duct on the DomainNet dataset. We consider a two-stage domain-incremental learning scenario, each containing five classes, and utilize t-SNE Van der Maaten & Hinton (2008) to visualize the representation space before and after Duct. We use dots to represent the classes in the first domain and triangles to represent classes in the second domain. We utilize the shadow regions to denote the decision boundary obtained by the classifier weights to represent different classes. As shown in Figure 4, although the embedding space before Duct shows competitive performance, there are two major flaws: 1) There exists the confusion region between yellow and red dots, indicating that previous classes are partially forgotten as data evolves. 2) The embedding space of the same class is still located in different regions, e.g., the purple and green classes, which is not ideal for domain-incremental learning. However, when applying Duct to consolidate the features and classifiers, the above problems are addressed in Figure 5. Specifically, since the consolidated embedding suits all tasks, the forgetting of previous domains is alleviated. Besides, the unified embedding space adequately places the same class of different domains together, favoring the final inference.
6 Conclusion
Domain-incremental learning is a desired ability for real-world learning systems to obtain new knowledge. In this paper, we propose dual consolidation (Duct) to achieve this goal. Since the forgetting in DIL occurs in two aspects, i.e., the embedding and the classifier, we separately design the consolidation technique to handle them. Specifically, we consider merging the pre-trained model with domain-specific task vectors to achieve the unified embedding. To compensate for the mismatch between consolidated features and classifiers, we design the classifier consolidation process by introducing semantic-guided transport. Extensive experiments validate the effectiveness of Duct.
Limitations: Possible limitations include the utilization of PTMs since feature consolidation relies on generalizable initialization. Future work includes extending Duct to non-PTM scenarios.
Appendix
In the main paper, we present a method to prevent forgetting in domain-incremental learning through representation and classifier consolidation. The supplementary material provides additional details on the experimental results mentioned in the main paper, along with extra empirical evaluations and discussions. The organization of the supplementary material is as follows:
  • Section A presents additional experimental results, including the running time comparison, detailed performance on different task orders, and results with other pre-trained weights.
  • Section B introduces the details about the datasets adopted in the main paper, including the number of tasks and images and differently-ordered sequences of tasks.
  • Section C introduces the compared methods adopted in the main paper.
Table 3: Average and last performance of different methods with the 1st task order in Section B.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implementations with exemplars (10 per class).
Method Office-Home DomainNet Core50 CDDB
$\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$
Finetune 73.48 76.23 47.99 35.80 74.44 72.18 50.43 50.23
Replay 80.62 83.80 64.68 62.88 85.71 91.66 50.26 50.91
iCaRL Rebuffi et al. (2017) 77.63 82.03 60.25 59.05 76.79 81.60 49.63 49.77
MEMO Zhou et al. (2023) 70.17 59.80 60.65 60.33 64.89 67.55 53.00 53.67
SimpleCIL Zhou et al. (2024a) 71.60 75.72 39.61 44.08 70.81 74.80 54.52 63.40
L2P Wang et al. (2022d) 74.49 78.44 49.00 51.07 83.47 88.33 58.69 67.89
DualPrompt Wang et al. (2022c) 75.88 80.72 51.92 52.73 85.42 88.09 63.10 71.72
CODA-Prompt Smith et al. (2023) 81.20 85.17 58.49 58.21 87.70 90.85 63.33 72.28
EASE Zhou et al. (2024b) 78.56 77.15 53.04 45.63 86.28 88.98 66.47 72.18
RanPAC McDonnell et al. (2023) 78.72 82.28 53.48 54.97 79.44 81.32 72.47 81.04
S-iPrompt Wang et al. (2022b) 77.32 80.42 58.21 58.88 80.91 82.09 61.75 73.23
Duct 83.76 86.79 67.28 68.52 92.06 94.83 78.32 84.98
Table 4: Average and last performance of different methods with the 2rd task order in Section B.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implementations with exemplars (10 per class).
Method Office-Home DomainNet Core50 CDDB
$\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$
Finetune 76.50 74.37 36.94 32.04 72.70 76.93 50.99 47.88
Replay 83.04 83.48 59.22 62.27 85.90 93.36 56.52 62.45
iCaRL Rebuffi et al. (2017) 81.48 81.90 53.75 57.91 75.54 81.17 61.08 76.31
MEMO Zhou et al. (2023) 71.44 62.92 53.41 57.04 61.30 67.52 50.54 48.59
SimpleCIL Zhou et al. (2024a) 68.45 75.72 38.18 44.08 71.63 74.80 63.19 63.40
L2P Wang et al. (2022d) 75.30 79.03 45.15 50.90 83.63 87.57 67.01 70.56
DualPrompt Wang et al. (2022c) 75.41 80.70 46.31 52.29 84.09 87.55 59.85 72.07
CODA-Prompt Smith et al. (2023) 81.21 84.53 52.86 57.38 87.91 91.21 63.05 73.53
EASE Zhou et al. (2024b) 75.96 74.76 47.31 44.62 86.32 86.69 64.08 69.19
RanPAC McDonnell et al. (2023) 78.03 82.28 50.11 54.98 79.44 81.32 75.00 81.04
S-iPrompt Wang et al. (2022b) 78.27 80.81 54.51 60.06 82.09 83.72 61.46 72.43
Duct 81.95 86.90 62.04 68.17 91.70 93.99 80.72 84.91
Table 5: Average and last performance of different methods with the 3rd task order in Section B.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implementations with exemplars (10 per class).
Method Office-Home DomainNet Core50 CDDB
$\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$
Finetune 81.82 74.93 44.84 29.45 77.35 79.12 51.68 52.81
Replay 86.98 84.10 67.23 60.52 84.91 91.60 87.52 75.30
iCaRL Rebuffi et al. (2017) 84.40 81.64 61.74 54.41 75.54 81.17 88.66 86.05
MEMO Zhou et al. (2023) 75.50 63.43 62.97 56.99 68.42 71.78 53.94 52.55
SimpleCIL Zhou et al. (2024a) 76.32 75.72 45.74 44.08 69.67 74.80 56.64 63.40
L2P Wang et al. (2022d) 81.94 79.57 52.16 45.05 84.21 88.24 67.87 67.87
DualPrompt Wang et al. (2022c) 82.45 80.79 54.01 49.28 84.91 88.39 66.26 71.19
CODA-Prompt Smith et al. (2023) 63.05 73.53 60.91 56.08 88.58 91.20 68.77 76.26
EASE Zhou et al. (2024b) 81.45 74.76 52.19 40.81 86.26 85.26 67.48 72.18
RanPAC McDonnell et al. (2023) 83.84 82.28 57.57 54.08 78.07 81.62 78.46 81.04
S-iPrompt Wang et al. (2022b) 83.40 80.68 64.02 61.22 82.65 84.20 69.30 72.99
Duct 87.31 86.94 70.08 67.06 91.97 94.78 88.84 85.84
Table 6: Average and last performance of different methods with the 4th task order in Section B.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implementations with exemplars (10 per class).
Method Office-Home DomainNet Core50 CDDB
$\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$
Finetune 82.10 77.09 25.82 16.65 75.97 75.21 53.80 50.42
Replay 86.38 82.67 65.35 60.26 85.89 92.23 89.91 77.27
iCaRL Rebuffi et al. (2017) 83.38 78.71 58.30 51.55 68.25 72.91 92.67 88.60
MEMO Zhou et al. (2023) 71.84 65.06 62.28 54.18 68.20 69.76 87.53 80.93
SimpleCIL Zhou et al. (2024a) 81.31 75.72 40.06 44.08 71.73 74.80 66.56 63.40
L2P Wang et al. (2022d) 85.49 81.71 48.59 45.47 83.35 87.01 77.38 61.50
DualPrompt Wang et al. (2022c) 84.74 81.02 50.90 46.77 85.23 86.89 81.46 72.89
CODA-Prompt Smith et al. (2023) 71.06 74.84 60.17 55.89 88.10 91.79 79.73 73.98
EASE Zhou et al. (2024b) 84.69 74.76 48.28 42.92 86.30 86.69 70.23 50.48
RanPAC McDonnell et al. (2023) 86.59 82.28 54.97 53.48 79.31 81.32 85.28 79.66
S-iPrompt Wang et al. (2022b) 85.52 80.19 61.32 60.89 82.11 83.50 81.28 73.06
Duct 89.79 86.98 64.12 64.70 92.12 94.56 89.38 84.31
Table 7: Average and last performance of different methods with the 5th task order in Section B.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implemented with exemplars (10 per class).
Method Office-Home DomainNet Core50 CDDB
$\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$
Finetune 77.69 78.18 38.49 26.9 76.73 77.52 53.40 49.21
Replay 84.12 84.72 67.45 59.88 85.37 92.22 50.34 50.12
iCaRL Rebuffi et al. (2017) 80.99 81.26 62.11 54.14 74.86 77.55 50.10 49.77
MEMO Zhou et al. (2023) 66.95 64.24 70.29 63.49 61.18 64.61 53.36 50.35
SimpleCIL Zhou et al. (2024a) 80.75 75.72 51.15 44.08 70.78 74.80 63.06 63.40
L2P Wang et al. (2022d) 81.37 81.38 57.33 51.13 83.21 88.22 65.68 54.42
DualPrompt Wang et al. (2022c) 82.55 81.00 57.33 48.12 82.99 83.48 71.00 69.18
CODA-Prompt Smith et al. (2023) 63.33 72.28 66.85 57.41 87.33 92.81 71.06 74.84
EASE Zhou et al. (2024b) 85.13 80.23 51.68 44.62 86.37 87.47 70.65 60.75
RanPAC McDonnell et al. (2023) 84.31 82.28 61.48 54.98 79.52 81.32 83.39 79.64
S-iPrompt Wang et al. (2022b) 82.70 80.47 67.73 61.23 81.98 83.38 68.76 72.09
Duct 88.55 86.92 72.29 66.59 91.88 94.21 83.46 85.48
Appendix A Supplied Experimental Results
In this section, we supply additional experiments to show the effectiveness of Duct, including the running time comparison, the detailed performance among different task orders, and more results with other pre-trained weights.
A.1 Running time comparison
In this section, we report the running time comparison among different compared methods. As shown in Figure 7, Duct costs competitive running time against other compared methods while having the best performance.
Figure 7: Running time comparison among different methods. Duct shows the best performance while having competitive training costs.
A.2 Performance of different task orders
In the main paper, we conduct experiments on the benchmark datasets with five task orders and report the average performance. In this section, we report the performance on each order in Table 3, 4, 5, 6, 7. The task sequences are reported in Section B.2.
Table 8: Details on domain size, train/test split, and instance number of the benchmark datasets. The dataset split and selection follows Wang et al. (2022c; d; b); Smith et al. (2023).
Domains Size Test set
CDDB-Hard biggan 4.0k standard splits
gaugan 10.0k 75%:25%
san 440 (’san’ - 80%:20%)
whichfaceisreal 2.0k
wild 10.5k
CORe50 s1 14.9k s3, s7, s10
s2 14.9k Indoor:Outdoor
s4 14.9k
s5 14.9k
s6 14.9k
s8 14.9k
s9 14.9k
s11 14.9k
DomainNet clipart 48.1k standard splits
infograph 51.6k 70%:30%
painting 72.2k
quickdraw 172.5k
real 172.9k
sketch 69.1k
Office-Home Art 2.4k random splits
Clipart 4.3k 70%:30%
Product 4.4k
Real World 4.3k
A.3 Different backbones
In the main paper, we mainly report the performance with ViT-B/16-IN1K. Since Duct is robust with the change of backbones, we report the performance with ViT-B/16-IN21K in this section. Apart from the benchmark datasets reported in Figure 2, the performance on the rest of the benchmarks is shown in Figure 8.
(a) Office-Home
(b) CDDB
Figure 8: Incremental performance of different methods with ViT-B/16 IN21K. We report the performance gap after the last incremental stage between Duct and the runner-up method at the end of the line.
Table 9: Task orders of CDDB-Hard.
CDDB-Hard Task 1 Task 2 Task 3 Task 4 Task 5
Order 1 san whichfaceisreal biggan wild gaugan
Order 2 wild whichfaceisreal san gaugan biggan
Order 3 biggan gaugan wild whichfaceisreal san
Order 4 gaugan biggan wild whichfaceisreal san
Order 5 whichfaceisreal san gaugan biggan wild
Table 10: Task orders of CORe50.
CORe50 Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Task 7 Task 8
Order 1 s11 s4 s2 s9 s1 s6 s5 s8
Order 2 s2 s9 s1 s6 s5 s8 s11 s4
Order 3 s4 s1 s9 s2 s5 s6 s8 s11
Order 4 s1 s9 s2 s5 s6 s8 s11 s4
Order 5 s9 s2 s5 s6 s8 s11 s4 s1
Table 11: Task orders of DomainNet.
DomainNet Task 1 Task 2 Task 3 Task 4 Task 5 Task 6
Order 1 clipart infograph painting quickdraw real sketch
Order 2 infograph painting quickdraw real sketch clipart
Order 3 painting quickdraw real sketch clipart infograph
Order 4 quickdraw real sketch clipart infograph painting
Order 5 real quickdraw painting sketch infograph clipart
Table 12: Task orders of Office-Home.
Office-Home Task 1 Task 2 Task 3 Task 4
Order 1 Art Clipart Product Real_World
Order 2 Clipart Product Real_World Art
Order 3 Product Clipart Real_World Art
Order 4 Real_World Product Clipart Art
Order 5 Art Real_World Product Clipart
Appendix B Dataset Details
In this section, we introduce the details about datasets, including the dataset information (i.e., the number of tasks and instances) and split information (i.e., the five splits adopted in the main paper).
B.1 Dataset introduction
We report the details about benchmark datasets in Table 8, and they are listed as follows.
  • DomainNet Peng et al. (2019) [^1] https://ai.bu.edu/M3SDA/ is a dataset of common objects in six different domains. All domains include 345 categories of objects, such as Bracelets, planes, birds, and cellos. The domains include clipart — a collection of clipart images; real — photos and real-world images; sketch — sketches of specific objects; infograph — infographic images with specific objects; painting — artistic depictions of objects in the form of paintings, and quickdraw — drawings of the worldwide players of the game ‘Quick Draw!’. We use the officially recommended version ‘Cleaned’ in this paper.
  • Office-Home Venkateswara et al. (2017) [^2] https://hemanthdv.github.io/officehome-dataset/ is a benchmark dataset for domain adaptation which contains four domains where each domain consists of 65 categories. The four domains are Art — artistic images in the form of sketches, paintings, ornamentation, etc.; Clipart — a collection of clipart images; Product — images of objects without a background; and Real-World — images of objects captured with a regular camera. It contains 15,500 images, with an average of around 70 images per class and a maximum of 99 images in a class.
  • CDDB-Hard Li et al. (2023) [^3] https://coral79.github.io/CDDB_web/. As a dataset mixed with real-world and model-generated images, the continual deepfake detection benchmark (CDDB) aims to simulate real-world deepfakes’ evolution. The authors put out three scenarios for evaluation, ‘EASY,’ ‘Hard,’ and ’Long,’ respectively. Following Wang et al. (2022b), we choose the ’Hard’ track for evaluation since it poses a more challenging problem due to its complexity.
  • CORe50 Lomonaco & Maltoni (2017) [^4] https://vlomonaco.github.io/core50/index.html#dataset. Composed of 11 sessions characterized by different backgrounds and lighting, CORe50 is built for continual object recognition. Numerous RGB-D images are divided into 8 indoor sessions for training and 3 outdoor sessions for testing, and each session includes a sequence of about 300 frames for all 50 objects.
B.2 Domain Sequences
In domain-incremental learning, different algorithms’ performances may be influenced by the order of domains. Consequently, we randomly shuffle the domains and organize five domain orders in the main paper, which are further utilized for a holistic evaluation. The task orders are reported in Table 9, 10, 11, 12.
Appendix C Compared Methods
In this section, we introduce the methods that were compared in the main paper. Note that we re-implement all methods using the same pre-trained model as initialization. They are listed as follows.
  • Finetune is a simple baseline in DIL, which directly optimizes the model with cross-entropy loss. It will suffer catastrophic forgetting since there is no restriction on preserving previous knowledge.
  • Replay Ratcliff (1990) is an exemplar-based method, which saves a set of exemplars from previous domains (i.e., in this paper, we save 10 exemplars per class) and replay them when learning new domains. Hence, forgetting can be alleviated since the model can revisit informative instances from previous domains when learning new ones. Of note, for classes with fewer than 10 instances, repeatable sampling is allowed to conform to the requirement.
  • iCaRL Rebuffi et al. (2017) is a knowledge distillation-based continual learning algorithm, which saves the previous model in memory. During updating, apart from the cross-entropy loss for learning new tasks, it also introduces the knowledge distillation loss between old and new models to avoid forgetting. It also requires saving an increasing number of exemplars.
  • MEMO Zhou et al. (2023) is an expansion-based continual learning algorithm that partially expands the network to catch new features. As for the implementation, we follow the original paper to decouple the network and expand the last transformer block for each new task.
  • SimpleCIL Zhou et al. (2024a) proposes this simple baseline in pre-trained model-based continual learning. It freezes the backbone representation, extracts the class center of each class, and utilizes a cosine classifier updated by assigning class centers to the classifier weights.
  • L2P Wang et al. (2022d) is the first work introducing prompt tuning in continual learning. With the pre-trained weights frozen, it learns a prompt pool containing many prompts. During training and inference, instance-specific prompts are selected to produce the instance-specific embeddings. However, as alluded to before, learning new domains will lead to the overwriting of existing prompts, thus triggering forgetting.
  • DualPrompt Wang et al. (2022c) extends L2P in two aspects. Apart from the prompt pool and prompt selection mechanism, it further introduces prompts instilled at different depths and task-specific prompts. During training and inference, the instance-specific and task-specific prompts work together to adjust the embeddings.
  • CODA-Prompt Smith et al. (2023) aims to avoid the prompt selection cost in L2P. It treats prompts in the prompt pool as bases and utilizes the attention results to combine multiple prompts as the instance-specific prompt.
  • EASE Zhou et al. (2024b) designs lightweight feature expansion technique with adapters to learn new features as data devolves. To fetch a classifier with the same dimension as ever-expanding features, it utilizes class-wise similarity to complete missing class prototypes.
  • RanPAC McDonnell et al. (2023) extends SimpleCIL by randomly projecting the features into the high-dimensional space and learning the online LDA classifier for final classification.
  • S-iPrompt Wang et al. (2022b) is specially designed for pre-trained model-based domain-incremental learning. It learns task-specific prompts for each domain and saves domain centers in the memory with K-Means. During inference, it first forwards the features to select the nearest domain center via KNN search. Afterward, the selected prompt will be appended to the input.

### v2

@@
 Extensive experimental results on four benchmark datasets demonstrate Duct’s state-of-the-art performance.
+Code is available at https://github.com/Estrella-fugaz/CVPR25-Duct.
+22footnotetext: Correspondence to: Han-Jia Ye (yehj@lamda.nju.edu.cn)
 1 Introduction
-Recent years have seen the rise of deep learning, demonstrating its strong potential in real-world applications He et al.
-(2016); Deng et al.
-(2009); Floridi & Chiriatti (2020).
-However, in a dynamic and ever-changing world, data often evolves in a stream format Aggarwal (2018); Ye et al.
-(2024), necessitating continuous updates with data from new domains Gama et al.
-(2014).
-For instance, autonomous vehicles must handle driving tasks across different seasons and weather conditions Bojarski et al.
-(2016), and face recognition systems are supposed to recognize users despite the varying lighting conditions Zhao et al.
-(2003).
-Correspondingly, Domain-Incremental Learning (DIL) van de Ven et al.
-(2022) addresses this challenge by absorbing new knowledge from new data distributions while preserving existing knowledge.
-However, sequentially updating the model with new domains often biases the embedding Yu et al.
-(2020) and classifier Zhao et al.
-(2020); Wu et al.
-(2019) modules toward the latest domain, leading to catastrophic forgetting Mermillod et al.
-(2013); De Lange et al.
-(2021) of previous knowledge.
+Recent years have seen the rise of deep learning, demonstrating its strong potential in real-world applications [22, 16, 18, 77, 43, 44].
+However, in a dynamic and ever-changing world, data often evolves in a stream format [1, 31], necessitating continuous updates with data from new domains [20].
+For instance, autonomous vehicles must handle driving tasks across different seasons and weather conditions [10], and face recognition systems are supposed to recognize users despite the varying lighting conditions [82].
+Correspondingly, Domain-Incremental Learning (DIL) [61] addresses this challenge by absorbing new knowledge from new data distributions while preserving existing knowledge.
+However, sequentially updating the model with new domains often biases the embedding [78] and classifier [81, 74] modules toward the latest domain, leading to catastrophic forgetting [41, 15] of previous knowledge.
 Therefore, developing methods for learning new domains without forgetting existing knowledge becomes a critical challenge for the machine learning community.
-To combat catastrophic forgetting, recent works leverage strong pre-trained models (PTMs)Han et al.
-(2021) as the initialization for DILWang et al.
-(2022d; c; b).
-PTMs’ robust representation abilities provide a powerful starting point, showing promising results in DIL benchmarks.
-These approaches freeze the pre-trained backbone to preserve existing knowledge and append lightweight modules Jia et al.
-(2022); Chen et al.
-(2022); Lian et al.
-(2022) to capture new patterns.
+To combat catastrophic forgetting, recent works leverage strong pre-trained models (PTMs) [21] as the initialization for DIL [70, 69, 68].
+PTMs’ robust feature representation abilities provide a powerful starting point, showing promising results in DIL benchmarks.
+These approaches freeze the pre-trained backbone to preserve existing knowledge and append lightweight modules [28, 14, 33] to capture new patterns.
 However, since incoming domains will also overwrite the appended modules, the representations still suffer from forgetting, leading to corresponding bias in the classifier.
@@
 Continually adjusting the features to capture the latest task biases the representations toward the newest domain.
-For example, suppose the previous task involves real-world photos while the subsequent task includes quick-draw images.
-The features will be overwhelmingly adjusted to highlight non-objective details, which are less helpful in distinguishing the previous domain, thus leading to forgetting.
+Assuming the previous task involves real-world photos while the subsequent one consists of quick-draw images, the features will be overwhelmingly adjusted to highlight non-objective details.
+This may hinder the model from distinguishing the previous domain, resulting in forgetting.
 In light of this, an ideal feature space in DIL should fit all domains and equally highlight the domain-specific features for all seen tasks.
 However, as features become less generalizable due to overwriting, the classifier also becomes biased toward the latest domain.
-Since the classifier plays the pivotal role of mapping embeddings to classes, using the biased classifier can cause detrimental effects to decision-making.
-Consequently, it is imperative to consolidate existing knowledge both in the feature and classifier to resist forgetting.
+Since the classifier plays the pivotal role of mapping embeddings to classes, using the biased classifier can cause detrimental effects on decision-making.
+Consequently, it is imperative to consolidate existing knowledge at both the representation and classifier levels to resist forgetting.
 There are two main challenges to achieving this goal.
@@
 This merged embedding consolidates task-specific information from all previous domains, providing informative features without forgetting.
-Consequently, we achieve non-forgettable features incrementally and resist feature-level forgetting.
+Consequently, we obtain non-forgettable features incrementally and resist feature-level forgetting.
 To map updated features to classes, we design a classifier consolidation process that merges historical classifier weights with calibrated weights.
 By extracting class-wise semantic relationships in the joint space, we develop a classifier transport mechanism that estimates classifiers of previous domain classes in the latest embedding space.
-Through consolidating features and classifiers in this coordinated way, Duct balances all seen domains and robustly resists forgetting, achieving competitive performance on various benchmarks.
+Through consolidating features and classifiers in this coordinated way, Duct balances all seen domains and robustly resists forgetting, achieving strong performance on various benchmarks.
 2 Related Work
-Continual Learning/Incremental Learning is a popular topic in the machine learning field, aiming to incorporate new knowledge within streaming data De Lange et al.
-(2021); Masana et al.
-(2023); Wang et al.
-(2024); Zhao et al.
-(2021); Zheng et al.
-(2024); Qi et al.
-(2024).
-It can be further classified into three subcategories, i.e., task-incremental learning (TIL) De Lange et al.
-(2021), class-incremental learning (CIL) Masana et al.
-(2023), and domain-incremental learning (DIL) van de Ven et al.
-(2022).
+Continual Learning/Incremental Learning aims to incorporate new knowledge within streaming data [15, 38, 67].
+It can be further classified into three subcategories, i.e., task-incremental learning (TIL) [15], class-incremental learning (CIL) [38, 87], and domain-incremental learning (DIL) [61].
 Among them, TIL and CIL are faced with emerging new classes and are required to learn them without forgetting.
-By contrast, DIL mainly focuses on the scenario where label space is fixed while incoming data involves new domains Wang et al.
-(2022b); Shi & Wang (2023).
-In continual learning, typical solutions include using data replay Bang et al.
-(2021); Aljundi et al.
-(2019b); Isele & Cosgun (2018); Rolnick et al.
-(2019); Ratcliff (1990) to review former knowledge, using knowledge distillation Hinton et al.
-(2015); Rebuffi et al.
-(2017); Li & Hoiem (2017) or parameter regularization Kirkpatrick et al.
-(2017); Zenke et al.
-(2017); Chaudhry et al.
-(2018); Aljundi et al.
-(2018; 2019a) to maintain existing knowledge.
-Recent advances also resort to bias correction Hou et al.
-(2019); Zhao et al.
-(2020); Ahn et al.
-(2021); Castro et al.
-(2018); Wu et al.
-(2019); Belouadah & Popescu (2019) to rectify the model bias or expand the network structure as data evolves Rusu et al.
-(2016); Aljundi et al.
-(2017); Yan et al.
-(2021); Wang et al.
-(2022a); Zhou et al.
-(2023).
-Domain-Incremental Learning with Pre-Trained Models: With the rapid development of pre-training techniques, PTMs provide a strong initialization for DIL models to boost the feature generalizability Smith et al.
-(2023); Wang et al.
-(2022c; d; b).
-Current PTM-based DIL methods mainly resort to visual prompt tuning Jia et al.
-(2022) to adjust the pre-trained features.
+By contrast, DIL mainly focuses on the scenario where label space is fixed while incoming data involves new domains [68, 56, 42].
+In continual learning, typical solutions include using data replay [8, 7, 27, 53, 51, 83, 34, 35] to review former knowledge, using knowledge distillation [23, 52, 32] or parameter regularization [29, 79, 13, 5, 6, 25] to maintain existing knowledge.
+Recent advances also resort to bias correction [24, 81, 2, 12, 74, 9] to rectify the model bias or expand the network structure as data evolves [55, 4, 75, 65, 85, 84, 66].
+Domain-Incremental Learning with PTMs: With the rapid development of pre-training techniques, PTMs provide a strong initialization for DIL models to boost the feature generalizability [58, 69, 70, 68, 60, 49, 37, 76].
+Current PTM-based DIL methods mainly resort to visual prompt tuning [28] to adjust the pre-trained features.
 With the pre-trained features frozen, it learns a prompt pool as external knowledge and selects instance-specific prompts to encode task information.
-L2P Wang et al.
-(2022d) utilizes the query-key matching mechanism to select instance-specific prompts, while DualPrompt Wang et al.
-(2022c) extends it by learning task-specific and instance-specific prompts jointly.
-By contrast, S-Prompts Wang et al.
-(2022b) learns domain-specific prompts for DIL and retrieves prompts via KNN search.
-To alleviate the prompt selection cost, CODA-Prompt Smith et al.
-(2023) modifies the prompt selection process into an attention-based weighted combination.
-Apart from prompt tuning, there are also some methods that directly build classifiers upon the pre-trained features Zhou et al.
-(2024a); McDonnell et al.
-(2023).
-Model Merging: Recent advances in pre-training have made weight interpolation among different models a useful technique in model editing Ainsworth et al.
-(2023); Frankle et al.
-(2020); Ilharco et al.
-(2023); Matena & Raffel (2021); Wortsman et al.
-(2022a; b).
+L2P [70] utilizes the query-key matching mechanism to select instance-specific prompts, while DualPrompt [69] extends it by learning task-specific and instance-specific prompts jointly.
+By contrast, S-Prompts [68] learns domain-specific prompts for DIL and retrieves prompts via KNN search.
+To alleviate the prompt selection cost, CODA-Prompt [58] reformulates this selection step as an attention-based weighted combination process.
+There are also some methods that directly build classifiers upon the pre-trained features [88, 40].
+Model Merging: Recent advances in pre-training have made weight interpolation among different models a useful technique in model editing [3, 19, 26, 39, 72, 73].
 Most work focuses on simple averaging of multiple models, aiming to enhance model in terms of its performance on the single task and robustness to out-of-distribution data.
-In this paper, we consider merging the pre-trained weights and task vector parameters to cultivate multitask representation for all seen domains.
+This paper considers merging pre-trained weights and task vectors to cultivate multitask representation for all seen domains.
 3 From Old Domains to New Domains
-In this section, we introduce the background information about domain-incremental learning and the baselines for applying pre-trained models to domain-incremental learning.
 3.1 Domain-Incremental Learning
-In domain-incremental learning, the model is faced with the data stream with a sequence of tasks $\left\{\mathcal{D}^{1},\mathcal{D}^{2},\cdots,\mathcal{D}^{B}\right\}$, where $\mathcal{D}^{b}=\left\{\mathcal{X}_{b},\mathcal{Y}_{b}\right\}$ is the $b$-th training task, i.e., $\mathcal{X}_{b}=\left\{{\bf x}_{i}\right\}_{i=1}^{n_{b}}$ and $\mathcal{Y}_{b}=\left\{y_{i}\right\}_{i=1}^{n_{b}}$.
+In domain-incremental learning, the model is faced with the data stream containing a sequence of tasks $\left\{\mathcal{D}^{1},\mathcal{D}^{2},\cdots,\mathcal{D}^{B}\right\}$, where $\mathcal{D}^{b}=\left\{\mathcal{X}_{b},\mathcal{Y}_{b}\right\}$ is the dataset of the $b$-th training task, i.e., $\mathcal{X}_{b}=\left\{{\bf x}_{i}\right\}_{i=1}^{n_{b}}$ and $\mathcal{Y}_{b}=\left\{y_{i}\right\}_{i=1}^{n_{b}}$.
 Each training instance ${\bf x}_{i}\in\mathbb{R}^{D}$ belongs to class $y_{i}\in Y$.
-This paper follows the exemplar-free setting Zhu et al.
-(2021); Wang et al.
-(2022d), i.e., during the $b$-th training stage, we can only access data in the current dataset $\mathcal{D}^{b}$.
+This paper follows the exemplar-free setting [89, 70], i.e., during the $b$-th training stage, we can only access data in the current dataset $\mathcal{D}^{b}$.
 In DIL, the label space $Y$ is unchanged throughout the learning process, while the distribution of input instance keeps changing from domain to domain, i.e., $p(\mathcal{X}_{b})\neq p(\mathcal{X}_{b^{\prime}})$ for $b\neq b^{\prime}$.
@@
 Consequently, DIL models are supposed to classify instances of new domains while not forgetting previous ones.
-Following Wang et al.
-(2022c; d; b); Smith et al.
-(2023), we assume a pre-trained Vision Transformer (ViT) Dosovitskiy et al.
-(2020) is available as the initialization for $f({\bf x})$.
+Following [69, 70, 68, 58], we assume a pre-trained Vision Transformer (ViT) [17] is available as the initialization for $f({\bf x})$.
 For simplicity, we decouple the network structure into the embedding function $\phi(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}^{d}$ (i.e., the final [CLS] token) and the linear classifier $W\in\mathbb{R}^{d\times|Y|}$.
@@
 Although the label space is static during the learning process, data distribution of the same class across different domains can yield significant domain gaps.
-Hence, we follow existing works Wang et al.
-(2022c; d; b); Smith et al.
-(2023) to expand the classifier $W\in\mathbb{R}^{d\times b|Y|}$ as new tasks emerge, and the final prediction is made by $\left(\operatorname*{arg\,max}_{i}{\bf w}_{i}^{\top}\phi({\bf x})\right)\textbf{ mod }|Y|$.
+Hence, we follow existing works [69, 70, 68, 58] to scale the classifier up to $W\in\mathbb{R}^{d\times b|Y|}$ when the $b$-th task emerges, and the final prediction is made by $\left(\mathop{\arg\max}_{i}{\bf w}_{i}^{\top}\phi({\bf x})\right)\textbf{ mod }|Y|$.
 3.2 Baselines in Domain-Incremental Learning
-In domain-incremental learning, a naive solution facing the incoming datasets of new domains is to directly optimize the embedding and classifier:
+In domain-incremental learning, a naive solution to handle the incoming datasets of new domains is to directly optimize the embedding and classifier:
 $\min_{W\cup\phi}\sum_{({\bf x},y)\in{D^{b}}}\ell\left(W^{\top}\phi({\bf x}),y\right)\,.$ (2)
 With pre-trained weights as initialization, Eq.
-2 can quickly capture the discriminative features within the new domain.
-However, since the embedding $\phi(\cdot)$ is kept changing among different domains, it quickly loses the generalization ability to previous domains and suffers from forgetting.
-To resist feature-level forgetting, a feasible solution is to freeze the pre-trained weights and append lightweight modules to encode domain-specific knowledge, e.g., prompt Jia et al.
-(2022).
-A representative work L2P Wang et al.
-(2022d) organizes a set of prompts as the prompt pool (i.e., $\mathbf{Pool}$) and optimizes the model via:
-$\min_{\mathbf{Pool}\cup W}\sum_{({\bf x},y)\in{D^{b}}}\ell\left(W^{\top}\bar{\phi}\left({\bf x};\mathbf{Pool}\right),y\right)+\mathcal{L}_{\mathbf{Pool}}\,,$ (3)
-where $\bar{\phi}\left({\bf x};\mathbf{Pool}\right)$ denotes the prompted feature representation with backbone frozen, and $\mathcal{L}_{\mathbf{Pool}}$ denotes the prompt selection loss Wang et al.
-(2022d).
+(2) can quickly capture the discriminative features within the new domain.
+However, since the embedding $\phi(\cdot)$ keeps changing across different domains, it quickly loses the generalization ability to previous domains and suffers from forgetting.
+To resist feature-level forgetting, a feasible solution is to freeze the pre-trained weights and append lightweight modules to encode domain-specific knowledge, e.g., prompt [28].
+A representative work L2P [70] organizes a set of prompts as the prompt pool ($\mathcal{P}$) and optimizes:
+$\min_{\mathcal{P}\cup W}\sum_{({\bf x},y)\in{D^{b}}}\ell\left(W^{\top}\bar{\phi}\left({\bf x};\mathcal{P}\right),y\right)+\mathcal{L}_{\mathcal{P}}\,,$ (3)
+where $\bar{\phi}\left({\bf x};\mathcal{P}\right)$ denotes the prompted feature representation with backbone frozen, and $\mathcal{L}_{\mathcal{P}}$ denotes the prompt selection loss [70].
 Eq.
-3 dynamically retrieves instance-specific prompts to adjust the representations and learns the classifier to map the features to corresponding classes.
+(3) dynamically retrieves instance-specific prompts to adjust the representations and learns the classifier to map the features to corresponding classes.
 Since the backbone weights are frozen, it can alleviate the representation drift during updating.
 Discussions: Although Eq.
-2 and Eq.
-3 adopt different policies for model updating, both of them suffer from the feature-level forgetting.
+(2) and Eq.
+(3) adopt different policies for model updating, both of them suffer from the feature-level forgetting.
 Specifically, tuning the model via Eq.
-2 can encode task-specific knowledge into the embedding, but also comes with the risk of feature being fully overwritten by new domains since there is no restriction on retaining previous knowledge.
+(2) can encode task-specific knowledge into the embedding, but also comes with the risk of features being fully overwritten by new domains since there is no restriction on retaining previous knowledge.
 Besides, although Eq.
-3 freezes the pre-trained weights, the trainable prompts appended may also incur forgetting.
+(3) freezes the pre-trained weights, the trainable prompts appended may also incur forgetting.
 Since the prompted feature relies on instance-specific prompt selection, optimizing the prompts for new domains still has the risk of overwriting previous ones.
-With the features being biased toward the latest domain, the classifier is also adjusted upon it.
+With the features being biased toward the latest domain, the classifier is adjusted accordingly.
 Consequently, the classifiers of previous tasks are incompatible with the unstable and ever-changing features, resulting in the classifier-level forgetting.
@@
 Since sequentially updating the backbone or prompts will both result in forgetting, a proper solution is needed to make full use of all historical features and consolidate them into a unified one.
-Secondly, as the representation changes from task to task, the mismatch between classifier and embedding becomes more severe as the data evolves.
+Secondly, as the embedding space changes from task to task, the mismatch between representation and classifier becomes more severe as the data evolves.
 Consequently, a classifier consolidation step is needed to calibrate them to be compatible with the embeddings.
-In the following sections, we introduce the learning paradigm for representation consolidation and classifier consolidation. Lastly, we provide detailed guidelines for training and inference.
 4.1 Representation Consolidation
@@
 Those models are obtained by optimizing the same PTM via Eq.
-2, thus having the discriminability for each specific domain and can be seen as the ‘domain expert.’ In an oversimplified scenario where we know which domain the instance is from, we can directly utilize the corresponding expert for prediction.
-However, since such auxiliary information is unavailable in DIL, we need to obtain a omnipotent embedding space that suits all domains.
+(2), thus having the discriminability for each specific domain and can be seen as the ‘domain expert.’ In an oversimplified scenario where we know which domain the instance is from, we can utilize the corresponding expert for prediction.
+However, since such auxiliary information is unavailable in DIL, we need to obtain an omnipotent embedding space that suits all domains.
 In this way, there is no need to decide which embedding to use since the omnipotent embedding is strong enough to capture all domain-specific features.
-Correspondingly, we take inspiration from the model merging community Ilharco et al.
-(2023); Ramé et al.
-(2023) that an ideal model for multiple domains can be achieved by combining multiple task vectors.
+Correspondingly, we take inspiration from model merging [26, 50] and cognitive-developmental theory [48] that an ideal model for multiple domains can be achieved by combining multiple task vectors.
 We denote the relative change of the embedding function as $\delta_{\phi_{i}}=\phi_{i}-\phi_{0}$, where $\phi_{0}$ is the weight of the pre-trained model.
 Hence, $\delta_{\phi_{i}}$ represents the relative change of the weights in the $i$-th domain, and we can build the unified embedding via:
-$\phi^{m}_{i}=\phi_{0}+\alpha_{\phi}\sum_{i}\delta_{\phi_{i}}\,,$ (4)
-which combines the pre-trained weight with every task vector.
-Since all the fine-tuned embeddings start from the same pre-trained weight, Ilharco et al.
-(2023); Zhang & Bottou (2023) verify the task vectors share low similarity due to the semantic gap between different domains.
+$\phi^{m}_{i}=\phi_{0}+\alpha_{\phi}\sum_{k=1}^{i}\delta_{\phi_{k}}\,,$ (4)
+where $\alpha_{\phi},\phi^{m}_{i}$ refers to the weight of task vector $\delta_{\phi_{k}}$ and the merged backbone respectively.
+For task vector start from the same pre-trained weight, [26, 80] verify the task vectors share low similarity due to the semantic gap between different domains.
 Consequently, adding such weights enables the merged model to highlight all task-specific features, and Eq.
-4 provides a feasible way to obtain the universal feature representation that suits all seen domains.
+(4) provides a feasible way to obtain the universal feature representation that suits all seen domains.
 Representation consolidation with task similarity: Although Eq.
-4 provides a simple way to consolidate multiple task vectors to build a unified embedding, it still lacks consideration in measuring task-wise similarities.
-For example, if two distinct domains are more similar, highlighting those features would be more helpful for recognizing classes of corresponding domains.
-Hence, we introduce task-similarity ($\textbf{Sim}_{0,i}$) into the representation consolidation process and replace Eq.
-4 into:
-$\phi^{m}_{i}=\phi_{0}+\alpha_{\phi}\sum_{i}\textbf{Sim}_{0,i}\delta_{\phi_{i}}\,,$ (5)
-Specifically, since the merged embedding is built upon pre-trained weights and task vectors, we consider the similarity of pre-trained model $\phi_{0}$ and subsequent models $\phi_{i}$ to adjust the merging process.
-A naive way to is to directly measure the cosine similarity between $\phi_{0}$ and $\phi_{i}$.
-However, since pre-trained weights are with millions of dimensions, the similarity in such space is ineffective due to curse of dimensionality.
-To this end, we utilize the current training task as an indicator for task similarity.
+(4) provides a simple way to consolidate multiple task vectors to build a unified embedding, it still lacks consideration in measuring task-wise similarities.
+For example, if two distinct domains are more similar, highlighting those features would be more helpful for recognizing corresponding domains.
+Hence, we introduce task-similarity ($\textbf{Sim}_{0,i}$) into the consolidation process.
+To measure the similarity, a naive way is to directly measure the similarity between the embedding function of pre-trained model $\phi_{0}$ and that of the subsequent model $\phi_{i}$.
+However, since the weights of pre-trained model are with millions of dimensions, the similarity in such space is ineffective due to the curse of dimensionality.
+To this end, we utilize the current training task as an indicator of task similarity.
 Specifically, for every class in $\mathcal{D}^{b}$, we utilize the backbone $\phi_{i}$ to extract class centers in its embedding space:
@@
 $\textbf{Sim}_{0,i}=\frac{1}{|{Y}|}\sum^{|{Y}|}_{j=1}\text{sim}({\bm{c}}^{0}_{j},{\bm{c}}^{i}_{j})\,,$ (6)
-where we utilize cosine similarity to calculate center-wise similarity $\text{sim}(\cdot,\cdot)$.
+where we utilize cosine similarity to calculate center-wise similarity $\text{sim}(\cdot,\cdot)$. We then introduce task similarity into Eq. (4), and the unified embedding is denoted as:
+$\phi^{m}_{i}=\phi_{0}+\alpha_{\phi}\sum_{k=1}^{i}\textbf{Sim}_{0,k}\delta_{\phi_{k}}\,,$ (7)
 Effect of representation consolidation: Figure 1 (top) visualizes the representation consolidation process, where the merged backbone can unify multiple domains.
 Through Eq.
-5, we are able to build the joint embedding space that suits all tasks by combining them in the parameter space.
+(7), we are able to build the joint embedding space that suits all tasks by combining them in the parameter space.
 Since tasks in DIL emerge one-by-one, the representation consolidation process can also be made incrementally, i.e., $\phi^{m}_{i}=\phi^{m}_{i-1}+\alpha_{\phi}\textbf{Sim}_{0,i}\delta_{\phi_{i}}$.
@@
 In each training stage, we first fine-tune the model via Eq.
-2 and conduct representation consolidation to aggregate the representations.
+(2) and then conduct representation consolidation to aggregate the representations.
 In this way, we obtain a unified embedding space across all seen tasks.
@@
 In Eq.
-5, we design the feature consolidation process, which aggregates multiple fine-tuned models by accumulating task vectors.
+(7), we design the representation consolidation process, which aggregates multiple fine-tuned models by accumulating task vectors.
 However, a fatal problem still exists, i.e., there is a mismatch between the classifiers and the consolidated embeddings.
-Since the classifiers are optimized to match the embeddings to the corresponding class, the matching degree drastically decays as the backbone is replaced with another one.
-Recalling the background information in Section 3.1 that we utilize an independent classifier for each domain, we need to design the extra classifier consolidation process to align the classifiers to the embedding space.
+Since the classifiers are optimized to align the embeddings with the corresponding class, the matching degree can decrease drastically when the embedding space is totally updated.
+Recalling the background information in Section 3.1 that we utilize an independent classifier for each domain, we need to design an extra classifier consolidation process to align the classifiers with the updated embedding space.
 In each training stage, we denote the classifier for previous domains as ‘old classifier’ ($W_{o}$) and the classifier for the current domain as ‘new classifier’ ($W_{n}$).
 We then discuss how to align them with the consolidated features.
-New Classifier Retraining: In each training stage, we have the training data $\mathcal{D}^{b}$ in hand.
-Consequently, it is intuitive to coordinate the new classifier with the consolidated features via:
+New Classifier Retraining: In each training stage, we have the training data $\mathcal{D}^{b}$ in hand. Hence, it is intuitive to coordinate the new classifier with the consolidated features via:
 $\min_{W_{n}}\sum_{({\bf x},y)\in{D^{b}}}\ell\left(W_{n}^{\top}\bar{\phi}_{i}^{m}({\bf x}),y\right)\,,$ (8)
-where the consolidated feature $\bar{\phi}_{i}^{m}(\cdot)$ is frozen to resist further mismatch, and we can align the new classifier ${W_{n}}$ to match the features with Eq. 8.
-Old Classifier Transport: Eq.
-8 aligns the new classifier to the latest embedding space.
-However, a fatal problem still exists, i.e., the old classifier is also incompatible with the merged embedding space.
+where the consolidated feature $\bar{\phi}_{i}^{m}(\cdot)$ is frozen to resist further mismatch, and we can adjust the new classifier ${W_{n}}$ to match the features using Eq. (8).
+Old Classifier Transport: Though Eq.
+(8) aligns the new classifier to the latest embedding space, another issue remains, i.e., the old classifier is also incompatible with the merged embedding space.
 Consequently, the previous knowledge of old domains shall be forgotten in the incremental learning process.
 If we have plenty of training instances of previous domains, a similar calibration process can be done as in Eq.
-8.
+(8).
 However, due to the exemplar-free restrictions, we cannot save any previous instances in the memory.
@@
 Implementing $\mathcal{T}$ via Optimal Transport: The function $\mathcal{T}$ encodes the correlation between the set of features among two tasks.
-Considering that the weight matrix of a linear layer reveals the relative relationship among feature-class pairs and the final logit is the aggregation of all features, we can design the recombination of exiting classifiers with a linear mapping $T\in\mathbb{R}^{\beta\times\gamma}$.
+Considering that the weight matrix of a linear layer reveals the relative relationship among feature-class pairs and the final logit is the aggregation of all features, we can design the recombination of existing classifiers with a linear mapping $T\in\mathbb{R}^{\beta\times\gamma}$.
 $T$ encodes the cross-task correlation between the current and the previous domains, where larger values in $T$ denote higher class-wise similarity.
@@
 We denote $\bm{\mu}_{1}\in\Delta_{\beta}$ and $\bm{\mu}_{2}\in\Delta_{\gamma}$, where $\Delta_{d}=\left\{\bm{\mu}:\bm{\mu}\in\mathbb{R}_{+}^{d},\bm{\mu}^{\top}\mathbf{1}=1\right\}$ is the $d$-dimensional simplex.
-$\bm{\mu}_{1}$ and $\bm{\mu}_{2}$ denote the importance of each class across domains, and we set them as uniform distribution.
-To map a set of classes to another between old and new domains, a cost matrix $Q\in\mathbb{R}_{+}^{\beta\times\gamma}$ is further introduced to guide the transition.
+$\bm{\mu}_{1}$ and $\bm{\mu}_{2}$ denote the importance of each class across domains, and we set them as uniform distributions.
+To map the distribution of $\beta$ new classes to that of $\gamma$ classes from old domains, a cost matrix $Q\in\mathbb{R}_{+}^{\beta\times\gamma}$ is further introduced to guide the transition.
 The larger weight of $Q_{i,j}$ indicates we need to pay more cost when reusing the classifier of $i$-th class to assist the $j$-th class.
-Consequently, the matrix $T$ can be formulated as the coupling of two distributions, aiming to connect classes from different domains at the lowest transportation cost.
+Consequently, the transport matrix $T$ can be formulated as the coupling of two distributions, aiming to connect classes from different domains at the lowest transportation cost.
 Hence, $T$ can be obtained via minimizing:
 $\min_{T}\langle T,Q\rangle\quad\text{ s.t. }T\mathbf{1}=\bm{\mu}_{1},T^{\top}\mathbf{1}=\bm{\mu}_{2},T\geq 0\,,$ (9)
-which is the Kantorovich formulation of optimal transport (OT) Villani (2008); Peyré et al.
-(2019).
+which is the Kantorovich formulation of optimal transport (OT) [64, 47].
 Optimizing Eq.
-9 enables us to show how to move the probability mass across domains with low cost.
+(9) enables us to show how to move the probability mass across domains with low cost.
 Since the target is to reuse the retrained classifier $W_{n}$ to adjust previous classifier $W_{o}$, we set $\bm{\mu}_{1}\in\Delta_{|Y|}$ and $\bm{\mu}_{2}\in\Delta_{|Y_{o}|}$ with uniform class marginals, where $|Y_{o}|$ denotes the number of classes in $W_{o}$.
 Solving Eq.
-9 produces the alignment between different domains, and we can apply $T$ to the new classifier to obtain the estimated old classifier, i.e., $\hat{W}_{o}=W_{n}T$.
+(9) establishes the mapping relationship between different domains, and we can apply $T$ to the new classifier to obtain the estimated old classifier, i.e., $\hat{W}_{o}=W_{n}T$.
 Defining the Transportation Cost with Semantic Information: Solving Eq.
-9 requires a proper definition of the cross-domain cost, i.e., $Q$.
+(9) requires a proper definition of the cross-domain cost, i.e., $Q$.
 The higher cost indicates it is less effective to transport the classifier to the target class and vice versa.
@@
 Specifically, before the training of each stage, we utilize the pre-trained backbone $\phi_{0}$ to extract class centers in its embedding space via Eq.
-6, i.e., $[{\bm{c}}^{0}_{1},{\bm{c}}^{0}_{2},\cdots,{\bm{c}}^{0}_{|{Y}|}]$.
+(5), i.e., $[{\bm{c}}^{0}_{1},{\bm{c}}^{0}_{2},\cdots,{\bm{c}}^{0}_{|{Y}|}]$.
 These class centers indicate the most representative embedding of the corresponding classes.
@@
 Consequently, we calculate the Euclidean distance between class centers as the transportation cost, i.e., $Q_{i,j}=\left\|{\bm{c}}^{0}_{i}-{\bm{c}}^{0}_{j}\right\|_{2}^{2}$.
-Here classes $i$ and $j$ are from different domains.
+Here, classes $i$ and $j$ are from different domains.
 With the pair-wise transportation cost, we are able to optimize Eq.
-9 for the transportation plan $T$.
+(9) for the transportation plan $T$.
 Afterward, we utilize the interpolation between old classifiers and the transported classifier for consolidation:
@@
 The first step is designed to align the new classifiers with the unified embedding, while the second step aims to recombine the old classifiers with the calibrated classifier.
-In this way, we obtain a compatible classifier with the consolidated features, thus favoring the recognition in the unified embedding space.
+In this way, we obtain a unified classifier compatible with the consolidated features, thus favoring the recognition in the unified embedding space.
+(a) CDDB ViT-B/16 IN1K
+(b) CORe50 ViT-B/16 IN1K
+(c) DomainNet ViT-B/16 IN1K
+Figure 2: Incremental performance of different methods with the same pre-trained model.
+We report the performance gap after the last incremental stage between Duct and the runner-up method at the end of the line.
+Table 1: Average and last performance of different methods among five task orders.
+The best performance is shown in bold.
+All methods are implemented with ViT-B/16 IN1K.
+Methods with $\dagger$ indicate implemented with exemplars (10 per class).
+Method Office-Home DomainNet CORe50 CDDB
+$\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$
+Finetune 78.32$\pm$3.28 76.16$\pm$1.39 28.17$\pm$6.47 38.82$\pm$7.65 75.44$\pm$1.68 76.19$\pm$2.36 52.08$\pm$1.35 50.11$\pm$1.62
+Replay [51] 84.23$\pm$2.31 83.75$\pm$0.68 64.78$\pm$2.98 61.16$\pm$1.19 85.56$\pm$0.38 92.21$\pm$0.63 66.91$\pm$18.0 63.21$\pm$11.6
+iCaRL [52] 81.66$\pm$2.43 81.11$\pm$1.23 59.89$\pm$2.86 57.46$\pm$2.31 74.43$\pm$3.18 79.86$\pm$2.96 68.43$\pm$18.7 70.50$\pm$16.5
+MEMO [85] 71.18$\pm$2.76 63.09$\pm$1.80 61.92$\pm$5.39 58.41$\pm$3.20 64.80$\pm$3.16 68.24$\pm$2.41 60.87$\pm$13.4 58.09$\pm$11.7
+SimpleCIL [88] 75.69$\pm$5.03 75.72$\pm$0.00 42.95$\pm$4.84 44.08$\pm$0.00 70.92$\pm$0.74 74.80$\pm$0.00 60.80$\pm$4.49 63.40$\pm$0.00
+L2P [70] 79.72$\pm$4.19 80.03$\pm$1.29 50.45$\pm$4.10 48.72$\pm$2.83 83.57$\pm$0.35 87.87$\pm$0.51 67.33$\pm$5.98 64.45$\pm$5.83
+DualPrompt [69] 80.20$\pm$3.81 80.85$\pm$0.14 52.28$\pm$3.35 50.46$\pm$3.17 84.53$\pm$0.89 87.27$\pm$1.06 68.33$\pm$7.52 71.41$\pm$1.24
+CODA-Prompt [58] 84.70$\pm$2.94 85.07$\pm$0.34 59.85$\pm$4.49 59.99$\pm$0.88 87.92$\pm$0.41 91.57$\pm$0.69 69.19$\pm$6.11 74.18$\pm$1.33
+EASE [86] 81.16$\pm$3.52 76.33$\pm$2.16 50.50$\pm$2.27 43.72$\pm$1.70 86.30$\pm$0.04 87.02$\pm$1.21 67.78$\pm$2.44 64.96$\pm$8.36
+RanPAC [40] 82.30$\pm$3.34 82.28$\pm$0.00 55.20$\pm$3.93 54.80$\pm$0.36 79.16$\pm$0.55 81.38$\pm$0.12 78.92$\pm$4.85 80.48$\pm$0.68
+S-iPrompt [68] 81.50$\pm$3.17 80.51$\pm$0.21 61.16$\pm$4.57 60.46$\pm$0.90 81.95$\pm$0.57 83.38$\pm$0.70 68.51$\pm$7.20 72.76$\pm$0.43
+Duct 86.27$\pm$2.95 86.91$\pm$0.06 67.16$\pm$3.75 67.01$\pm$1.35 91.95$\pm$0.15 94.47$\pm$0.33 84.14$\pm$4.37 85.10$\pm$0.52
 4.3 Summary of Duct
+In Duct, facing a new training task, we extract the class centers with $\phi_{0}$.
+Afterward, we optimize the model and separately consolidate the features and classifiers.
+The consolidated model ($[W_{o}^{m};W_{n}]^{\top}\phi^{m}_{i}$) is then utilized for testing.
+We utilize Sinkhorn’s algorithm [57] to solve the OT problem in Eq.
+(9).
+We summarize the training steps with pseudo code in the supplementary material.
+Discussions on memory cost: During the training process, we need to keep two models in the memory, i.e., the historical consolidated backbone $\phi^{m}_{i-1}$ and the online backbone $\phi_{i}$ for learning the current task.
+However, after the learning process of each domain, the current running model is merged into $\phi^{m}_{i}$ (Eq.
+(7)) without the need to be stored in memory.
+During inference, we utilize the consolidated backbone $\phi^{m}_{i}$ and classifiers $[W_{o}^{m};W_{n}]$ for classification, which is the same as a single backbone.
+Consequently, Duct can be fairly compared with others.
+5 Experiment
+(a) Multiple trails
+(b) Forgetting measure
+(c) Parameter robustness
+Figure 3: Further analysis on multiple task orders, forgetting measure, and parameter robustness.
+(a): Incremental performance of different methods on CORe50 with five task orders.
+The shadow indicates standard deviation.
+(b): Forgetting measure (lower is better) of different methods on CDDB dataset among five task orders.
+Duct shows the least forgetting among all methods.
+(c): Average incremental performance with change of the consolidation ratios.
+In this section, we conduct experiments on benchmark datasets to compare Duct to existing state-of-the-art methods.
+We also provide ablation studies on the components and parameter robustness to analyze the effect of different modules.
+Visualizations of the embedding space also verify the effectiveness of our proposed method.
+5.1 Experimental Setup
+Dataset: Following the benchmark settings [69, 70, 68, 58] in PTM-based domain-incremental learning, we evaluate the performance on Office-Home [63], DomainNet [46], CORe50 [36], and CDDB-Hard [30].
+Specifically, Office-Home has four domains, DomainNet has six domains, CORe50 has 11 domains, and CDDB-Hard has five domains.
+We consider five task orders to shuffle these domains in the DIL setting for a holistic evaluation and report the details in the supplementary material.
+Comparison methods: In the comparison, we consider two types of methods, including exemplar-based methods, i.e., Replay [51], iCaRL [52], MEMO [85], and SOTA exemplar-free DIL methods, i.e., SimpleCIL [88], L2P [70], DualPrompt [69], CODA-Prompt [58], EASE [86], RanPAC [40], and S-iPrompt [68].
+We utilize the same pre-trained backbone for all compared methods.
+Implementation details: We deploy the experiments using PyTorch [45] and Pilot [59] on NVIDIA 4090.
+Following [70, 86], we consider two typical pre-trained weights, i.e., ViT-B/16-IN21K and ViT-B/16-IN1K.
+Both are pre-trained with ImageNet21K [54], while the latter is further finetuned on ImageNet1K.
+We optimize Duct using SGD optimizer with a batch size of 128 for 15 epochs.
+The learning rate is set to $0.001$.
+We select 10 exemplars per class for exemplar-based methods using herding [71] algorithm.
+In Duct, we set the consolidation parameter $\alpha_{\phi}=0.5,\alpha_{w}=0.5$.
+The code will be made publicly available upon acceptance.
+Performance Measure: Following [70, 68], we denote the accuracy among all seen domains after learning the $b$-th task as $\mathcal{A}_{b}$, we mainly consider $\mathcal{A}_{B}$ (the performance after learning the last stage) and $\bar{\mathcal{A}}=\frac{1}{B}\sum_{b=1}^{B}\mathcal{A}_{b}$ (the average performance among all incremental stages) for comparison.
+We also consider the forgetting measure [13] to measure the relative forgetting degree in DIL.
+5.2 Benchmark Comparison
+We report the results on benchmark datasets in Figure 2 and Table 1, with the remaining results provided in the supplementary material.
+As we can infer from these results, Duct consistently outperforms other compared methods by $\bf 1\sim 7$in the final accuracy.
+It must be noted that the differences of the starting point are due to the different tuning techniques, which cannot be directly aligned since the number of trainable parameters are different in those methods.
+Specifically, sequentially finetuning the model suffers from catastrophic forgetting, even starting with the pre-trained model.
+To compensate for the forgetting phenomena, we observe that several exemplar-based methods (Replay, iCaRL, and MEMO) show stable improvements to the baseline.
+However, since some of these works are designed for the class-incremental learning scenario, the expansion or distillation target may not be optimal for the current setting.
+We then compare Duct to the methods specially designed for PTMs and find prompt-based methods (L2P, DualPrompt, CODA-Prompt, and S-iPrompt) yield inferior performance to ours.
+We also observe that Duct is compatible with various pre-trained weights and shows stable improvements with ViT-B/16 IN1K and IN21K (see supplementary).
+Besides, we also report the incremental performance among five task orders (average and standard deviation) in Table 1 and Figure 3(a).
+As we can infer from the table, Duct works robustly across different task orders, showing stable improvements against other state-of-the-art methods.
+Variations $\bar{\mathcal{A}}$ ${\mathcal{A}}_{B}$
+Baseline 66.56 63.40
+Variation 1 80.36 74.17
+Variation 2 81.41 76.82
+Variation 3 85.42 80.31
+Duct 87.74 82.35
+Table 2: Ablation study on differnt modules in Duct.
+Figure 4: Before Duct.
+Figure 5: After Duct.
+Figure 6: Visualizations of the embedding space via t-SNE [62] on DomainNet.
+We visualize two incremental stages and utilize dots to represent the first domain and triangles for the second domain.
+The shadow region denotes the decision boundaries.
+5.3 Further Analysis
+Forgetting Measure: Figure 2 and Table 1 mainly focus on the accuracy measure, which utilizes all seen domains to measure the relative performance.
+Apart from the accuracy measure, we also follow [68, 70] to utilize the forgetting measure, which captures the stability of existing knowledge among previous domains.
+As shown in Figure 3(b), we report the forgetting measure among all compared methods on the CDDB dataset.
+As we can infer from the figure, several methods suffer from severe forgetting, e.g., Finetune and MEMO, indicating that they are unsuitable for the domain-incremental learning scenario.
+We also observe that prompt-based methods freeze the backbone to resist feature drift, which tackles the forgetting phenomena.
+However, Duct still shows the least forgetting (i.e., 0.12) among all compared methods, indicating its strong performance.
+Parameter Robustness: There are two major consolidation steps in Duct, i.e., the representation consolidation in Eq.
+(7) and the classifier consolidation in Eq.
+(10).
+These consolidation steps involve the hyper-parameters for merging a set of models/classifiers, i.e., the backbone merge ratio $\alpha_{\phi}$ and the classifier merge ratio $\alpha_{W}$.
+In this section, we conduct ablations on the choice of these parameters to investigate the robustness of Duct to their changes.
+Specifically, we choose $\alpha_{\phi}$ and $\alpha_{W}$ among $\{0.1,0.25,0.5,0.75,1.0\}$, resulting in 25 parameter combinations.
+We conduct experiments with these parameter combinations on the Office-Home dataset and present the average performance in Figure 3(c).
+As we can infer from the figure, Duct is generally robust to parameter changes.
+Besides, since the merge parameters are designed to trade-off between old and new knowledge, we find either a small value (e.g., 0.1) or a large value (e.g., 1.0) works poorly.
+The poor performance of these marginal parameters also indicates the importance of the merging process since a small value or large value equals ignoring part of the important information during consolidation.
+By contrast, both of them prefer a mild value, i.e., choosing them around {0.25, 0.5} shows better performance.
+Hence, we suggest setting $\alpha_{\phi}=0.5,\alpha_{w}=0.5$ as the default setting.
+Compositional Components: In this section, we conduct experiments to investigate the importance of each module in Duct on CDDB dataset.
+As shown in Table 2, ‘Baseline’ denotes directly freezing the embedding and extracting class centers as the classifier, which shows poor performance due to the domain gap to the downstream domains.
+We then utilize Eq.
+(4) to sequentially merge the backbone to consolidate the representations and denote it as Variation 1.
+It shows that adding the representation consolidation process drastically improves the performance, indicating the superiority of such unified representation in DIL.
+However, when switching Eq.
+(4) to Eq.
+(7), we find Variation 2 further improves the performance, indicating that task similarity is helpful in obtaining a universal embedding space.
+Afterward, we combine it with the classifier retraining process in Eq.
+(8), which is denoted as Variation 3.
+It shows that the mismatch between consolidated features and the classifiers weakens the performance, and aligning the new classifiers to the embedding helps DIL.
+When comparing Duct to Variation 3, we find the old classifier transport is also vital to recover former knowledge and resist forgetting, which improves the final accuracy by 2%.
+Ablation studies verify the efficacy of different modules in Duct.
+Visualizations: In this section, we visualize the embedding space to show the effectiveness of Duct on the DomainNet dataset.
+We consider a two-stage domain-incremental learning scenario, each containing five classes, and utilize t-SNE [62] to visualize the representation space before and after Duct.
+We use dots to represent the classes in the first domain and triangles for classes in the second domain.
+We utilize the shadow regions to denote the decision boundary obtained by the classifier weights to represent different classes.
+As shown in Figure 4, although the embedding space before Duct shows competitive performance, there are two major flaws: 1) There exists the confusion region between yellow and red dots, indicating that previous classes are partially forgotten as data evolves.
+2) The embedding space of the same class is still located in different regions, e.g., the purple and green areas, which is not ideal for domain-incremental learning.
+However, when applying Duct to consolidate the features and classifiers, the above problems are addressed in Figure 5.
+Specifically, since the consolidated embedding suits all tasks, the forgetting of previous domains is alleviated.
+Besides, the unified embedding space adequately clusters the same class of different domains together, favoring the final inference.
+6 Conclusion
+Domain-incremental learning is a desired ability for real-world learning systems to obtain new knowledge.
+In this paper, we propose dual consolidation (Duct) to achieve this goal.
+Since the forgetting in DIL occurs in two aspects, i.e., the representation and the classifier, we separately design the consolidation technique to handle them.
+Specifically, we consider continually merging the pre-trained model with domain-specific task vectors to achieve the unified embedding.
+To compensate for the mismatch between consolidated features and classifiers, we design the classifier consolidation process by introducing semantic-guided transport.
+Extensive experiments validate the effectiveness of Duct.
+Limitations: Possible limitations include the utilization of PTMs since representation consolidation relies on generalizable initialization. Future work includes extending Duct to non-PTM scenarios.
+Acknowledgments.
+This work is partially supported by NSFC (U23A20382, 62476123, 62376118, 62006112, 62250069), Fundamental Research Funds for the Central Universities (2024300373,14380021), CCF-Tencent Rhino-Bird Open Research Fund RAGR20240101, Collaborative Innovation Center of Novel Software Technology and Industrialization.
+Appendix
+In the main paper, we present a method to prevent forgetting in domain-incremental learning through representation and classifier consolidation.
+The supplementary material provides additional details on the algorithm and experimental results mentioned in the main paper, along with extra empirical evaluations and discussions.
+The organization of the supplementary material is as follows:
+  • Section I provides the pseudocode of Duct.
+  • Section II presents detailed experimental results, including the running time comparison, detailed performance on different task orders, and results with other pre-trained weights.
+  • Section III introduces the details about the datasets adopted in the main paper, including the number of tasks and images and differently-ordered sequences of tasks.
+  • Section IV introduces the compared methods adopted in the main paper.
+  • Section V presents a series of additional experiments to further validate the robustness of Duct.
+Appendix I Algorithm Pipeline
+Following the notation presented in the main body, we summarize the training steps in Algorithm 1.
+First of all, we prepare the dataset and load the pre-trained model weights, then calculate the mean representation for each incoming class, as shown in Lines 2$\sim$3.
+Next, we fine-tune the model and extract the task vector, which is then used to apply the model merging technique, as described in Line 5.
+After that, we keep the merged backbone frozen and retrain the classification layer to align it with the consolidated representation.
+Finally, we apply classifier transport to the classifiers to transfer semantic information, as outlined in Lines 7$\sim$8.
 Algorithm 1 Duct for DIL
@@
 8:     Consolidate old classifiers via Eq. 10; return the updated model;
-We summarize the training steps in Algorithm 1.
-Specifically, facing a new training task, we extract the class centers with $\phi_{0}$.
-Afterward, we optimize the model and separately consolidate the features and classifiers.
-The consolidated model ($[W_{o}^{m};W_{n}]^{\top}\phi^{m}_{i}$) is then utilized for testing.
-We utilize Sinkhorn’s algorithm Sinkhorn & Knopp (1967) to solve the OT problem in Eq.
-9.
-Discussions on memory cost: During the training process, we need to keep two models in the memory, i.e., the historical consolidated backbone $\phi^{m}_{i}$ and the online backbone $\phi_{i}$ for learning the current task.
-However, after the learning process of each domain, the current running model is merged into $\phi^{m}_{i}$ (Eq.
-5) without needing to be kept in memory.
-During inference, we utilize the consolidated backbone $\phi^{m}_{i}$ and classifiers $[W_{o}^{m};W_{n}]$ for classification, which is the same as a single backbone.
-Consequently, Duct can be fairly compared with others.
-5 Experiment
-In this section, we conduct experiments on benchmark datasets to compare Duct to existing state-of-the-art methods.
-We also provide ablation studies on the components and parameter robustness to analyze the effect of different modules.
-Visualizations of the embedding space also verify the effectiveness of our proposed method.
-(a) Office-Home ViT-B/16 IN1K
-(b) DomainNet ViT-B/16 IN1K
-(c) CORe50 ViT-B/16 IN1K
-(d) CDDB ViT-B/16 IN1K
-(e) DomainNet ViT-B/16 IN21K
-(f) CORe50 ViT-B/16 IN21K
-Figure 2: Incremental performance of different methods with the same pre-trained model.
-We report the performance gap after the last incremental stage between Duct and the runner-up method at the end of the line.
-Table 1: Average and last performance of different methods among five task orders.
-The best performance is shown in bold.
-All methods are implemented with ViT-B/16 IN1K.
-Methods with $\dagger$ indicate implemented with exemplars (10 per class).
-Method Office-Home DomainNet CORe50 CDDB
-$\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$
-Finetune 78.32$\pm$3.28 76.16$\pm$1.39 28.17$\pm$6.47 38.82$\pm$7.65 75.44$\pm$1.68 76.19$\pm$2.36 52.08$\pm$1.35 50.11$\pm$1.62
-Replay Ratcliff (1990) 84.23$\pm$2.31 83.75$\pm$0.68 64.78$\pm$2.98 61.16$\pm$1.19 85.56$\pm$0.38 92.21$\pm$0.63 66.91$\pm$18.0 63.21$\pm$11.6
-iCaRL Rebuffi et al. (2017) 81.66$\pm$2.43 81.11$\pm$1.23 59.89$\pm$2.86 57.46$\pm$2.31 74.43$\pm$3.18 79.86$\pm$2.96 68.43$\pm$18.7 70.50$\pm$16.5
-MEMO Zhou et al. (2023) 71.18$\pm$2.76 63.09$\pm$1.80 61.92$\pm$5.39 58.41$\pm$3.20 64.80$\pm$3.16 68.24$\pm$2.41 60.87$\pm$13.4 58.09$\pm$11.7
-SimpleCIL Zhou et al. (2024a) 75.69$\pm$5.03 75.72$\pm$0.00 42.95$\pm$4.84 44.08$\pm$0.00 70.92$\pm$0.74 74.80$\pm$0.00 60.80$\pm$4.49 63.40$\pm$0.00
-L2P Wang et al. (2022d) 79.72$\pm$4.19 80.03$\pm$1.29 50.45$\pm$4.10 48.72$\pm$2.83 83.57$\pm$0.35 87.87$\pm$0.51 67.33$\pm$5.98 64.45$\pm$5.83
-DualPrompt Wang et al. (2022c) 80.20$\pm$3.81 80.85$\pm$0.14 52.28$\pm$3.35 50.46$\pm$3.17 84.53$\pm$0.89 87.27$\pm$1.06 68.33$\pm$7.52 71.41$\pm$1.24
-CODA-Prompt Smith et al. (2023) 84.70$\pm$2.94 85.07$\pm$0.34 59.85$\pm$4.49 59.99$\pm$0.88 87.92$\pm$0.41 91.57$\pm$0.69 69.19$\pm$6.11 74.18$\pm$1.33
-EASE Zhou et al. (2024b) 81.16$\pm$3.52 76.33$\pm$2.16 50.50$\pm$2.27 43.72$\pm$1.70 86.30$\pm$0.04 87.02$\pm$1.21 67.78$\pm$2.44 64.96$\pm$8.36
-RanPAC McDonnell et al. (2023) 82.30$\pm$3.34 82.28$\pm$0.00 55.20$\pm$3.93 54.80$\pm$0.36 79.16$\pm$0.55 81.38$\pm$0.12 78.92$\pm$4.85 80.48$\pm$0.68
-S-iPrompt Wang et al. (2022b) 81.50$\pm$3.17 80.51$\pm$0.21 61.16$\pm$4.57 60.46$\pm$0.90 81.95$\pm$0.57 83.38$\pm$0.70 68.51$\pm$7.20 72.76$\pm$0.43
-Duct 86.27$\pm$2.95 86.91$\pm$0.06 67.16$\pm$3.75 67.01$\pm$1.35 91.95$\pm$0.15 94.47$\pm$0.33 84.14$\pm$4.37 85.10$\pm$0.52
-5.1 Experimental Setup
-Dataset: Following the benchmark settings Wang et al.
-(2022c; d; b); Smith et al.
-(2023) in PTM-based domain-incremental learning, we evaluate the performance on Office-Home Venkateswara et al.
-(2017), DomainNet Peng et al.
-(2019), CORe50 Lomonaco & Maltoni (2017), and CDDB-Hard Li et al.
-(2023).
-Specifically, Office-Home has four domains, DomainNet has six domains, CORe50 has 11 domains, and CDDB-Hard has five domains.
-We consider five task orders to shuffle these domains in the DIL setting for a holistic evaluation and report the details in Section B.2.
-Comparison methods: In the comparison, we consider two types of methods, including exemplar-based methods, i.e., Replay Ratcliff (1990), iCaRL Rebuffi et al.
-(2017), MEMO Zhou et al.
-(2023), and SOTA exemplar-free DIL methods, i.e., SimpleCIL Zhou et al.
-(2024a), L2P Wang et al.
-(2022d), DualPrompt Wang et al.
-(2022c), CODA-Prompt Smith et al.
-(2023), EASE Zhou et al.
-(2024b), RanPAC McDonnell et al.
-(2023), and S-iPrompt Wang et al.
-(2022b).
-We utilize the same pre-trained backbone for all compared methods.
-Implementation details: We deploy the experiments using PyTorch Paszke et al.
-(2019) and Pilot Sun et al.
-(2023) on NVIDIA 4090.
-Following Wang et al.
-(2022d); Zhou et al.
-(2024b), we consider two typical pre-trained weights, i.e., ViT-B/16-IN21K and ViT-B/16-IN1K.
-Both are pre-trained with ImageNet21K Russakovsky et al.
-(2015), while the latter is further finetuned on ImageNet1K.
-We optimize Duct using SGD optimizer with a batch size of 128 for 15 epochs.
-The learning rate is set to $0.001$.
-We select 10 exemplars per class for exemplar-based methods using herding Welling (2009) algorithm.
-In Duct, we set the consolidation parameter $\alpha_{\phi}=0.5,\alpha_{w}=0.5$.
-The code will be made publicly available upon acceptance.
-Performance Measure: Following Wang et al.
-(2022d; b), we denote the accuracy among all seen domains after learning the $b$-th task as $\mathcal{A}_{b}$, we mainly consider $\mathcal{A}_{B}$ (the performance after learning the last stage) and $\bar{\mathcal{A}}=\frac{1}{B}\sum_{b=1}^{B}\mathcal{A}_{b}$ (the average performance among all incremental stages) for comparison.
-We also consider the forgetting measure Chaudhry et al.
-(2018) to measure the relative forgetting degree in DIL.
-5.2 Benchmark Comparison
-[…]
-Besides, we also report the incremental performance among five task orders (average and standard deviation) in Table 1 and Figure 3(a).
-As we can infer from the table, Duct works robustly among different task orders, showing stable improvements against other state-of-the-art methods.
-(a) Multiple trails
-(b) Forgetting measure $\downarrow$
-(c) Parameter robustness
-Figure 3: Further analysis on multiple task orders, forgetting measure, and parameter robustness.
-(a): Incremental performance of different methods on CORe50 with five task orders.
-The shadow indicates standard deviation.
-(b): Forgetting measure (lower is better) of different methods on CDDB dataset among five task orders.
-Duct shows the least forgetting among all methods.
-(c): Average incremental performance with change of the consolidation ratios.
-5.3 Further Analysis
-Forgetting Measure: Figure 2 and Table 1 mainly focus on the accuracy measure, which utilizes all seen domains to measure the relative performance.
-Apart from the accuracy measure, we also follow Wang et al.
-(2022b; d) to utilize the forgetting measure, which captures the stability of existing knowledge among previous domains.
-As shown in Figure 3(b), we report the forgetting measure among all compared methods on the CDDB dataset.
-As we can infer from the figure, several methods suffer from severe forgetting, e.g., Finetune and MEMO, indicating that they are unsuitable for the domain-incremental learning scenario.
-We also observe that prompt-based methods freeze the backbone to resist feature drift, which tackles the forgetting phenomena.
-However, Duct still shows the least forgetting (i.e., 0.12) among all compared methods, indicating its strong performance.
-Parameter Robustness: There are two major consolidation steps in Duct, i.e., the representation consolidation in Eq.
-5 and the classifier consolidation in Eq.
-10.
-These consolidation steps include the merging parameters between a set of models/classifiers, i.e., the backbone merge ratio $\alpha_{\phi}$ and the classifier merge ratio $\alpha_{W}$.
-In this section, we conduct ablations on the choice of these parameters to investigate the robustness of Duct with change of them.
-Specifically, we choose $\alpha_{\phi}$ and $\alpha_{W}$ among $\{0.1,0.25,0.5,0.75,1.0\}$, resulting in 25 parameter combinations.
-We conduct experiments with these parameter combinations on the Office-Home dataset and show the average performance in Figure 3(c).
-As we can infer from the figure, Duct is generally robust to parameter changes.
-Besides, since the merge parameters are designed to trade-off between old and new knowledge, we find a small value (e.g., 0.1) or large value (e.g., 1.0) works poorly.
-The poor performance of these marginal parameters also indicates the importance of the merging process since a small value or large value equals to ignoring part of the important information during consolidation.
-By contrast, both of them prefer a mild value, i.e., choosing them around {0.25, 0.5} shows better performance.
-Hence, we suggest setting $\alpha_{\phi}=0.5,\alpha_{w}=0.5$ as the default setting.
-Variations $\bar{\mathcal{A}}$ ${\mathcal{A}}_{B}$
-Baseline 66.56 63.40
-Variation 1 80.36 74.17
-Variation 2 81.41 76.82
-Variation 3 85.42 80.31
-Duct 87.74 82.35
-Table 2: Ablation study on differnt modules in Duct.
-Figure 4: Before Duct.
-Figure 5: After Duct.
-Figure 6: Visualizations of the embedding space via t-SNE Van der Maaten & Hinton (2008) on DomainNet.
-We visualize two incremental stages and utilize dots to represent the first domain and triangles for the second domain.
-The shadow region denotes the decision boundaries.
-Compositional Components: In this section, we conduct experiments to investigate the importance of each module in Duct on CDDB dataset.
-As shown in Table 2, ‘Baseline’ denotes directly freezing the embedding and extracting class centers as the classifier, which shows poor performance due to the domain gap to the downstream domains.
-We then utilize Eq.
-4 to sequentially merge the backbone to consolidate the representations and denote it as Variation 1.
-It shows that adding the representation consolidation process drastically improves the performance, indicating the superiority of such unified representation in DIL.
-However, when switching Eq.
-4 to Eq.
-5, we find Variation 2 further improves the performance, indicating that task similarity is helpful in obtaining a universal embedding space.
-Afterward, we combine it with the classifier retraining process in Eq.
-8, which is denoted as Variation 3.
-It shows that the mismatch between consolidated features and the classifiers weakens the performance, and aligning the new classifiers to the embedding helps DIL.
-When comparing Duct to Variation 3, we find the old classifier transport is also vital to recover former knowledge and resist forgetting, which improves the final accuracy by 2%.
-Ablation studies verify the efficacy of different modules in Duct.
-Visualizations: In this section, we visualize the embedding space to show the effectiveness of Duct on the DomainNet dataset.
-We consider a two-stage domain-incremental learning scenario, each containing five classes, and utilize t-SNE Van der Maaten & Hinton (2008) to visualize the representation space before and after Duct.
-We use dots to represent the classes in the first domain and triangles to represent classes in the second domain.
-We utilize the shadow regions to denote the decision boundary obtained by the classifier weights to represent different classes.
-As shown in Figure 4, although the embedding space before Duct shows competitive performance, there are two major flaws: 1) There exists the confusion region between yellow and red dots, indicating that previous classes are partially forgotten as data evolves.
-2) The embedding space of the same class is still located in different regions, e.g., the purple and green classes, which is not ideal for domain-incremental learning.
-However, when applying Duct to consolidate the features and classifiers, the above problems are addressed in Figure 5.
-Specifically, since the consolidated embedding suits all tasks, the forgetting of previous domains is alleviated.
-Besides, the unified embedding space adequately places the same class of different domains together, favoring the final inference.
-6 Conclusion
-Domain-incremental learning is a desired ability for real-world learning systems to obtain new knowledge.
-In this paper, we propose dual consolidation (Duct) to achieve this goal.
-Since the forgetting in DIL occurs in two aspects, i.e., the embedding and the classifier, we separately design the consolidation technique to handle them.
-Specifically, we consider merging the pre-trained model with domain-specific task vectors to achieve the unified embedding.
-To compensate for the mismatch between consolidated features and classifiers, we design the classifier consolidation process by introducing semantic-guided transport.
-Extensive experiments validate the effectiveness of Duct.
-Limitations: Possible limitations include the utilization of PTMs since feature consolidation relies on generalizable initialization. Future work includes extending Duct to non-PTM scenarios.
-Appendix
-In the main paper, we present a method to prevent forgetting in domain-incremental learning through representation and classifier consolidation.
-The supplementary material provides additional details on the experimental results mentioned in the main paper, along with extra empirical evaluations and discussions.
-The organization of the supplementary material is as follows:
-  • Section A presents additional experimental results, including the running time comparison, detailed performance on different task orders, and results with other pre-trained weights.
-  • Section B introduces the details about the datasets adopted in the main paper, including the number of tasks and images and differently-ordered sequences of tasks.
-  • Section C introduces the compared methods adopted in the main paper.
-Table 3: Average and last performance of different methods with the 1st task order in Section B.2.
+Appendix II Supplied Experimental Results
+In this section, we supply additional experiments to show the effectiveness of Duct, including the running time comparison, the detailed performance among different task orders, and more results with other pre-trained weights.
+II.1 Performance of different task orders
+In the main paper, we conduct experiments on the benchmark datasets with five task orders and report the average performance.
+In this section, we report the performance on each order in Table 3, 4, 5, 6, 7.
+The task sequences are reported in Section III.2.
+Table 3: Average and last performance of different methods with the 1st task order in Section III.2.
 The best performance is shown in bold.
@@
 $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$ $\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$
-Finetune 73.48 76.23 47.99 35.80 74.44 72.18 50.43 50.23
+Finetune 73.48 76.23 47.99 35.80 74.44 72.18 50.23 50.23
 Replay 80.62 83.80 64.68 62.88 85.71 91.66 50.26 50.91
-iCaRL Rebuffi et al. (2017) 77.63 82.03 60.25 59.05 76.79 81.60 49.63 49.77
-MEMO Zhou et al. (2023) 70.17 59.80 60.65 60.33 64.89 67.55 53.00 53.67
-SimpleCIL Zhou et al. (2024a) 71.60 75.72 39.61 44.08 70.81 74.80 54.52 63.40
-L2P Wang et al. (2022d) 74.49 78.44 49.00 51.07 83.47 88.33 58.69 67.89
-DualPrompt Wang et al. (2022c) 75.88 80.72 51.92 52.73 85.42 88.09 63.10 71.72
-CODA-Prompt Smith et al. (2023) 81.20 85.17 58.49 58.21 87.70 90.85 63.33 72.28
-EASE Zhou et al. (2024b) 78.56 77.15 53.04 45.63 86.28 88.98 66.47 72.18
-RanPAC McDonnell et al. (2023) 78.72 82.28 53.48 54.97 79.44 81.32 72.47 81.04
-S-iPrompt Wang et al. (2022b) 77.32 80.42 58.21 58.88 80.91 82.09 61.75 73.23
+iCaRL [52] 77.63 82.03 60.25 59.05 76.79 81.60 49.63 49.77
+MEMO [85] 70.17 59.80 60.65 60.33 64.89 67.55 53.00 53.67
+SimpleCIL [88] 71.60 75.72 39.61 44.08 70.81 74.80 54.52 63.40
+L2P [70] 74.49 78.44 49.00 51.07 83.47 88.33 58.69 67.89
+DualPrompt [69] 75.88 80.72 51.92 52.73 85.42 88.09 63.10 71.72
+CODA-Prompt [58] 81.20 85.17 58.49 58.21 87.70 90.85 63.33 72.28
+EASE [86] 78.56 77.15 53.04 45.63 86.28 88.98 66.47 72.18
+RanPAC [40] 78.72 82.28 53.48 54.97 79.44 81.32 72.47 81.04
+S-iPrompt [68] 77.32 80.42 58.21 58.88 80.91 82.09 61.75 73.23
 Duct 83.76 86.79 67.28 68.52 92.06 94.83 78.32 84.98
-Table 4: Average and last performance of different methods with the 2rd task order in Section B.2.
+Table 4: Average and last performance of different methods with the 2rd task order in Section III.2.
 The best performance is shown in bold.
@@
 Replay 83.04 83.48 59.22 62.27 85.90 93.36 56.52 62.45
-iCaRL Rebuffi et al. (2017) 81.48 81.90 53.75 57.91 75.54 81.17 61.08 76.31
-MEMO Zhou et al. (2023) 71.44 62.92 53.41 57.04 61.30 67.52 50.54 48.59
-SimpleCIL Zhou et al. (2024a) 68.45 75.72 38.18 44.08 71.63 74.80 63.19 63.40
-L2P Wang et al. (2022d) 75.30 79.03 45.15 50.90 83.63 87.57 67.01 70.56
-DualPrompt Wang et al. (2022c) 75.41 80.70 46.31 52.29 84.09 87.55 59.85 72.07
-CODA-Prompt Smith et al. (2023) 81.21 84.53 52.86 57.38 87.91 91.21 63.05 73.53
-EASE Zhou et al. (2024b) 75.96 74.76 47.31 44.62 86.32 86.69 64.08 69.19
-RanPAC McDonnell et al. (2023) 78.03 82.28 50.11 54.98 79.44 81.32 75.00 81.04
-S-iPrompt Wang et al. (2022b) 78.27 80.81 54.51 60.06 82.09 83.72 61.46 72.43
+iCaRL [52] 81.48 81.90 53.75 57.91 75.54 81.17 61.08 76.31
+MEMO [85] 71.44 62.92 53.41 57.04 61.30 67.52 50.54 48.59
+SimpleCIL [88] 68.45 75.72 38.18 44.08 71.63 74.80 63.19 63.40
+L2P [70] 75.30 79.03 45.15 50.90 83.63 87.57 67.01 70.56
+DualPrompt [69] 75.41 80.70 46.31 52.29 84.09 87.55 59.85 72.07
+CODA-Prompt [58] 81.21 84.53 52.86 57.38 87.91 91.21 63.05 73.53
+EASE [86] 75.96 74.76 47.31 44.62 86.32 86.69 64.08 69.19
+RanPAC [40] 78.03 82.28 50.11 54.98 79.44 81.32 75.00 81.04
+S-iPrompt [68] 78.27 80.81 54.51 60.06 82.09 83.72 61.46 72.23
 Duct 81.95 86.90 62.04 68.17 91.70 93.99 80.72 84.91
-Table 5: Average and last performance of different methods with the 3rd task order in Section B.2.
+Table 5: Average and last performance of different methods with the 3rd task order in Section III.2.
 The best performance is shown in bold.
@@
 Replay 86.98 84.10 67.23 60.52 84.91 91.60 87.52 75.30
-iCaRL Rebuffi et al. (2017) 84.40 81.64 61.74 54.41 75.54 81.17 88.66 86.05
-MEMO Zhou et al. (2023) 75.50 63.43 62.97 56.99 68.42 71.78 53.94 52.55
-SimpleCIL Zhou et al. (2024a) 76.32 75.72 45.74 44.08 69.67 74.80 56.64 63.40
-L2P Wang et al. (2022d) 81.94 79.57 52.16 45.05 84.21 88.24 67.87 67.87
-DualPrompt Wang et al. (2022c) 82.45 80.79 54.01 49.28 84.91 88.39 66.26 71.19
-CODA-Prompt Smith et al. (2023) 63.05 73.53 60.91 56.08 88.58 91.20 68.77 76.26
-EASE Zhou et al. (2024b) 81.45 74.76 52.19 40.81 86.26 85.26 67.48 72.18
-RanPAC McDonnell et al. (2023) 83.84 82.28 57.57 54.08 78.07 81.62 78.46 81.04
-S-iPrompt Wang et al. (2022b) 83.40 80.68 64.02 61.22 82.65 84.20 69.30 72.99
+iCaRL [52] 84.40 81.64 61.74 54.41 75.54 81.17 88.66 86.05
+MEMO [85] 75.50 63.23 62.97 56.99 68.42 71.78 53.94 52.55
+SimpleCIL [88] 76.32 75.72 45.74 44.08 69.67 74.80 56.64 63.40
+L2P [70] 81.94 79.57 52.16 45.05 84.21 88.24 67.87 67.87
+DualPrompt [69] 82.45 80.79 54.01 49.28 84.91 88.39 66.26 71.19
+CODA-Prompt [58] 63.05 73.53 60.91 56.08 88.58 91.20 68.77 76.26
+EASE [86] 81.45 74.76 52.19 40.81 86.26 85.26 67.48 72.18
+RanPAC [40] 83.84 82.28 57.57 54.08 78.07 81.62 78.46 81.04
+S-iPrompt [68] 83.40 80.68 64.02 61.22 82.65 84.20 69.30 72.99
 Duct 87.31 86.94 70.08 67.06 91.97 94.78 88.84 85.84
-Table 6: Average and last performance of different methods with the 4th task order in Section B.2.
+Table 6: Average and last performance of different methods with the 4th task order in Section III.2.
 The best performance is shown in bold.
@@
 Replay 86.38 82.67 65.35 60.26 85.89 92.23 89.91 77.27
-iCaRL Rebuffi et al. (2017) 83.38 78.71 58.30 51.55 68.25 72.91 92.67 88.60
-MEMO Zhou et al. (2023) 71.84 65.06 62.28 54.18 68.20 69.76 87.53 80.93
-SimpleCIL Zhou et al. (2024a) 81.31 75.72 40.06 44.08 71.73 74.80 66.56 63.40
-L2P Wang et al. (2022d) 85.49 81.71 48.59 45.47 83.35 87.01 77.38 61.50
-DualPrompt Wang et al. (2022c) 84.74 81.02 50.90 46.77 85.23 86.89 81.46 72.89
-CODA-Prompt Smith et al. (2023) 71.06 74.84 60.17 55.89 88.10 91.79 79.73 73.98
-EASE Zhou et al. (2024b) 84.69 74.76 48.28 42.92 86.30 86.69 70.23 50.48
-RanPAC McDonnell et al. (2023) 86.59 82.28 54.97 53.48 79.31 81.32 85.28 79.66
-S-iPrompt Wang et al. (2022b) 85.52 80.19 61.32 60.89 82.11 83.50 81.28 73.06
+iCaRL [52] 83.38 78.71 58.30 51.55 68.25 72.91 92.67 88.60
+MEMO [85] 71.84 65.06 62.28 54.18 68.20 69.76 87.53 80.93
+SimpleCIL [88] 81.31 75.72 40.06 44.08 71.73 74.80 66.56 63.40
+L2P [70] 85.49 81.71 48.59 45.47 83.35 87.01 77.38 61.50
+DualPrompt [69] 84.74 81.02 50.90 46.77 85.23 86.89 81.46 72.89
+CODA-Prompt [58] 71.06 74.84 60.17 55.89 88.10 91.79 79.73 73.98
+EASE [86] 84.69 74.76 48.28 42.92 86.30 86.69 70.23 50.48
+RanPAC [40] 86.59 82.28 54.97 53.48 79.31 81.32 85.28 79.66
+S-iPrompt [68] 85.52 80.19 61.32 60.89 82.11 83.50 81.28 73.06
 Duct 89.79 86.98 64.12 64.70 92.12 94.56 89.38 84.31
-Table 7: Average and last performance of different methods with the 5th task order in Section B.2.
+Table 7: Average and last performance of different methods with the 5th task order in Section III.2.
 The best performance is shown in bold.
@@
 Replay 84.12 84.72 67.45 59.88 85.37 92.22 50.34 50.12
-iCaRL Rebuffi et al. (2017) 80.99 81.26 62.11 54.14 74.86 77.55 50.10 49.77
-MEMO Zhou et al. (2023) 66.95 64.24 70.29 63.49 61.18 64.61 53.36 50.35
-SimpleCIL Zhou et al. (2024a) 80.75 75.72 51.15 44.08 70.78 74.80 63.06 63.40
-L2P Wang et al. (2022d) 81.37 81.38 57.33 51.13 83.21 88.22 65.68 54.42
-DualPrompt Wang et al. (2022c) 82.55 81.00 57.33 48.12 82.99 83.48 71.00 69.18
-CODA-Prompt Smith et al. (2023) 63.33 72.28 66.85 57.41 87.33 92.81 71.06 74.84
-EASE Zhou et al. (2024b) 85.13 80.23 51.68 44.62 86.37 87.47 70.65 60.75
-RanPAC McDonnell et al. (2023) 84.31 82.28 61.48 54.98 79.52 81.32 83.39 79.64
-S-iPrompt Wang et al. (2022b) 82.70 80.47 67.73 61.23 81.98 83.38 68.76 72.09
+iCaRL [52] 80.99 81.26 62.11 54.14 74.86 77.55 50.10 49.77
+MEMO [85] 66.95 64.24 70.29 63.49 61.18 64.61 53.36 50.35
+SimpleCIL [88] 80.75 75.72 51.15 44.08 70.78 74.80 63.06 63.40
+L2P [70] 81.37 81.38 57.33 51.13 83.21 88.22 65.68 54.42
+DualPrompt [69] 82.55 81.00 57.33 48.12 82.99 83.48 71.00 69.18
+CODA-Prompt [58] 63.33 72.28 66.85 57.41 87.33 92.81 71.06 74.84
+EASE [86] 85.13 80.23 51.68 44.62 86.37 87.47 70.65 60.75
+RanPAC [40] 84.31 82.28 61.48 54.98 79.52 81.32 83.39 79.64
+S-iPrompt [68] 82.70 80.47 67.73 61.23 81.98 83.38 68.76 72.09
 Duct 88.55 86.92 72.29 66.59 91.88 94.21 83.46 85.48
-Appendix A Supplied Experimental Results
-In this section, we supply additional experiments to show the effectiveness of Duct, including the running time comparison, the detailed performance among different task orders, and more results with other pre-trained weights.
-A.1 Running time comparison
-In this section, we report the running time comparison among different compared methods.
-As shown in Figure 7, Duct costs competitive running time against other compared methods while having the best performance.
-Figure 7: Running time comparison among different methods. Duct shows the best performance while having competitive training costs.
-A.2 Performance of different task orders
-In the main paper, we conduct experiments on the benchmark datasets with five task orders and report the average performance.
-In this section, we report the performance on each order in Table 3, 4, 5, 6, 7.
-The task sequences are reported in Section B.2.
-Table 8: Details on domain size, train/test split, and instance number of the benchmark datasets. The dataset split and selection follows Wang et al. (2022c; d; b); Smith et al. (2023).
+Table 8: Details on domain size, train/test split, and instance number of the benchmark datasets. The dataset split and selection follows [69, 70, 68, 58].
 Domains Size Test set
@@
 CORe50 s1 14.9k s3, s7, s10
-s2 14.9k Indoor:Outdoor
+s2 14.9k Indoor:Outdoor 8:3
 s4 14.9k
@@
 Real World 4.3k
-A.3 Different backbones
-In the main paper, we mainly report the performance with ViT-B/16-IN1K.
-Since Duct is robust with the change of backbones, we report the performance with ViT-B/16-IN21K in this section.
-Apart from the benchmark datasets reported in Figure 2, the performance on the rest of the benchmarks is shown in Figure 8.
-(a) Office-Home
-(b) CDDB
-Figure 8: Incremental performance of different methods with ViT-B/16 IN21K. We report the performance gap after the last incremental stage between Duct and the runner-up method at the end of the line.
 Table 9: Task orders of CDDB-Hard.
@@
 Order 5 Art Real_World Product Clipart
-Appendix B Dataset Details
+II.2 Different backbones
+In the main paper, we mainly report the performance using ViT-B/16-IN1K.
+In this section, we present performance on the remaining benchmark in Figure 7 and all the experimental results on ViT-B/16-IN21K in Figure 8, demonstrating the model’s robustness to changes in backbones.
+II.3 Running time comparison
+In this section, we report the running time comparison among different compared methods.
+As shown in Figure 9, Duct costs competitive running time against other compared methods while having the best performance.
+Figure 9: Running time comparison among different methods. Duct shows the best performance while having competitive training costs.
+Appendix III Dataset Details
 In this section, we introduce the details about datasets, including the dataset information (i.e., the number of tasks and instances) and split information (i.e., the five splits adopted in the main paper).
-B.1 Dataset introduction
+III.1 Dataset introduction
 We report the details about benchmark datasets in Table 8, and they are listed as follows.
-  • DomainNet Peng et al.
-(2019) [^1] https://ai.bu.edu/M3SDA/ is a dataset of common objects in six different domains.
+  • DomainNet [46]***https://ai.bu.edu/M3SDA/ is a dataset of common objects in six different domains.
 All domains include 345 categories of objects, such as bracelets, planes, birds, and cellos.
@@
 We use the officially recommended version ‘Cleaned’ in this paper.
-  • Office-Home Venkateswara et al.
-(2017) [^2] https://hemanthdv.github.io/officehome-dataset/ is a benchmark dataset for domain adaptation which contains four domains where each domain consists of 65 categories.
-The four domains are Art — artistic images in the form of sketches, paintings, ornamentation, etc.; Clipart — a collection of clipart images; Product — images of objects without a background; and Real-World — images of objects captured with a regular camera.
+  • Office-Home [63]https://hemanthdv.github.io/officehome-dataset/ is a benchmark dataset for domain adaptation which contains four domains where each domain consists of 65 categories.
+The four domains are Art — artistic images in the form of sketches, paintings, ornamentation, etc.; Clipart — a collection of clipart images; Product — images of objects without a background; and Real_World — images of objects captured with a regular camera.
 It contains 15,500 images, with an average of around 70 images per class and a maximum of 99 images in a class.
-  • CDDB-Hard Li et al.
-(2023) [^3] https://coral79.github.io/CDDB_web/.
+  • CDDB-Hard [30]https://coral79.github.io/CDDB_web/.
 As a dataset mixed with real-world and model-generated images, the continual deepfake detection benchmark (CDDB) aims to simulate real-world deepfakes’ evolution.
-The authors put out three scenarios for evaluation, ‘EASY,’ ‘Hard,’ and ’Long,’ respectively.
-Following Wang et al.
-(2022b), we choose the ’Hard’ track for evaluation since it poses a more challenging problem due to its complexity.
-  • CORe50 Lomonaco & Maltoni (2017) [^4] https://vlomonaco.github.io/core50/index.html#dataset.
+The authors put out three tracks for evaluation, ‘EASY,’ ‘Hard,’ and ’Long,’ respectively.
+Following [68], we choose the ’Hard’ track for evaluation since it poses a more challenging problem due to its complexity.
+  • CORe50 [36]§§§https://vlomonaco.github.io/core50/index.html#dataset.
 Composed of 11 sessions characterized by different backgrounds and lighting, CORe50 is built for continual object recognition.
 Numerous RGB-D images are divided into 8 indoor sessions for training and 3 outdoor sessions for testing, and each session includes a sequence of about 300 frames for all 50 objects.
-B.2 Domain Sequences
+III.2 Domain Sequences
 In domain-incremental learning, different algorithms’ performances may be influenced by the order of domains.
@@
 The task orders are reported in Table 9, 10, 11, 12.
-Appendix C Compared Methods
+Table 13: Incremental performances of methods on five representative pre-trained models, i.e., ResNet101, ViT-S/16, ViT-B/16-IN1K, ViT-B/16-IN21K, and ViT-B/16-DINO, comparing the selected best-performing algorithms and Duct.
+‘NA’ indicates that these methods are incompatible with ResNet.
+Duct achieves the optimal performances across all backbones.
+ResNet101 ViT-S/16 ViT-B/16-DINO ViT-B/16-IN1K ViT-B/16-IN21K
+CODA-Prompt [58] NA 52.90 55.47 58.49 58.99
+EASE [86] NA 40.84 38.72 53.04 53.11
+S-iPrompt [68] NA 55.12 60.12 58.21 60.13
+Duct 50.22 64.55 62.89 67.28 67.37
+Table 14: The upper bound on the performance of ViT-B/16-IN21K for all benchmarks, i.e., the performance achieved by fine-tuning on the union of the entire dataset.
+CDDB-Hard Core50 DomainNet Office-Home
+Upperbound 93.44 96.29 76.28 87.92
+Appendix IV Compared Methods
 In this section, we introduce the methods that were compared in the main paper. Note that we re-implement all methods using the same pre-trained model as initialization. They are listed as follows.
@@
 It will suffer catastrophic forgetting since there is no restriction on preserving previous knowledge.
-  • Replay Ratcliff (1990) is an exemplar-based method, which saves a set of exemplars from previous domains (i.e., in this paper, we save 10 exemplars per class) and replay them when learning new domains.
+  • Replay [51] is an exemplar-based method, which saves a set of exemplars from previous domains (i.e., in this paper, we save 10 exemplars per class) and replay them when learning new domains.
 Hence, forgetting can be alleviated since the model can revisit informative instances from previous domains when learning new ones.
 Of note, for classes with fewer than 10 instances, repeatable sampling is allowed to conform to the requirement.
-  • iCaRL Rebuffi et al.
-(2017) is a knowledge distillation-based continual learning algorithm, which saves the previous model in memory.
+  • iCaRL [52] is a knowledge distillation-based continual learning algorithm, which saves the previous model in memory.
 During updating, apart from the cross-entropy loss for learning new tasks, it also introduces the knowledge distillation loss between old and new models to avoid forgetting.
 It also requires saving an increasing number of exemplars.
-  • MEMO Zhou et al.
-(2023) is an expansion-based continual learning algorithm that partially expands the network to catch new features.
+  • MEMO [85] is an expansion-based continual learning algorithm that partially expands the network to catch new features.
 As for the implementation, we follow the original paper to decouple the network and expand the last transformer block for each new task.
-  • SimpleCIL Zhou et al.
-(2024a) proposes this simple baseline in pre-trained model-based continual learning.
+  • SimpleCIL [88] proposes this simple baseline in pre-trained model-based continual learning.
 It freezes the backbone representation, extracts the class center of each class, and utilizes a cosine classifier updated by assigning class centers to the classifier weights.
-  • L2P Wang et al.
-(2022d) is the first work introducing prompt tuning in continual learning.
+  • L2P [70] is the first work introducing prompt tuning in continual learning.
 With the pre-trained weights frozen, it learns a prompt pool containing many prompts.
@@
 However, as alluded to before, learning new domains will lead to the overwriting of existing prompts, thus triggering forgetting.
-  • DualPrompt Wang et al.
-(2022c) extends L2P in two aspects.
+  • DualPrompt [69] extends L2P in two aspects.
 Apart from the prompt pool and prompt selection mechanism, it further introduces prompts instilled at different depths and task-specific prompts.
 During training and inference, the instance-specific and task-specific prompts work together to adjust the embeddings.
-  • CODA-Prompt Smith et al.
-(2023) aims to avoid the prompt selection cost in L2P.
+  • CODA-Prompt [58] aims to avoid the prompt selection cost in L2P.
 It treats prompts in the prompt pool as bases and utilizes the attention results to combine multiple prompts as the instance-specific prompt.
-  • EASE Zhou et al.
-(2024b) designs lightweight feature expansion technique with adapters to learn new features as data devolves.
+  • EASE [86] designs lightweight feature expansion technique with adapters to learn new features as data devolves.
 To fetch a classifier with the same dimension as ever-expanding features, it utilizes class-wise similarity to complete missing class prototypes.
-  • RanPAC McDonnell et al. (2023) extends SimpleCIL by randomly projecting the features into the high-dimensional space and learning the online LDA classifier for final classification.
-  • S-iPrompt Wang et al.
-(2022b) is specially designed for pre-trained model-based domain-incremental learning.
+  • RanPAC [40] extends SimpleCIL by randomly projecting the features into the high-dimensional space and learning the online LDA classifier for final classification.
+  • S-iPrompt [68] is specially designed for pre-trained model-based domain-incremental learning.
 It learns task-specific prompts for each domain and saves domain centers in the memory with K-Means.
@@
 Afterward, the selected prompt will be appended to the input.
+Appendix V More Experiments
+To further demonstrate the effectiveness of Duct in different scenarios, we conduct a series of more comprehensive experiments in this section.
+V.1 More backbones
+Apart from ViT-B/16-IN1K and ViT-B/16-IN21K, we also evaluate additional backbones at different scales, including ResNet101 [22], ViT-S/16, and ViT-B/16-DINO [11].
+The results are reported in Table 13.
+Note that the compared methods are all based on prompt tuning technique [28] which is designed for Vision Transformers [17].
+Hence, these compared methods are incompatible with ResNet structures, while Duct can still boost ResNet in continual learning tasks, further highlighting that Duct is a general and versatile framework.
+The outcomes demonstrate that Duct outperforms all competitors across all PTMs.
+V.2 Upper bound performance
+To investigate the upper-bound performance of each dataset, we conduct joint training and present the results in Table 14.
+We adopt a learning rate of 0.0001 for the representation layer and 0.01 for the classification layer, with a batch size of 128.
+We report the optimal results in the table.
+V.3 Extending Duct to Class-Incremental Learning
+Since Duct is a general framework, it can also be applied to the class-incremental learning (CIL) scenario.
+By equally splitting 345 classes of DomainNet into 5 incremental stages, we formulate a class-incremental learning setting, each containing 69 classes.
+We keep the other settings the same as in the main paper and report the results in Table 15.
+The results indicate that DUCT can also be applied to the class-incremental learning setting, which we shall explore in future work.
+Table 15: Incremental performances of selected methods on CIL setting, evaluated on the most challenging benchmark dataset, DomainNet, using ViT-B/16-IN21K as the backbone.
+‘B0’ indicates that classes are uniformly split across tasks.
+The results indicate that Duct performs effectively in CIL scenarios.
+Method DomainNet B0 Inc69
+$\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$
+CODA-Prompt [58] 72.50 71.05
+EASE [86] 69.29 68.45
+S-iPrompt [68] 73.79 71.81
+Duct 75.29 74.68